### v2

# envname-P envname#1

Abstract
Prevailing code generation benchmarks, such as HumanEval+ and MBPP+, primarily evaluate large language models (LLMs) with pass@k on functional correctness using well-formed inputs. However, they ignore a crucial aspect of real-world software: adherence to contracts—the preconditions and validity constraints that dictate how ill-formed inputs must be rejected. This critical oversight means that existing benchmarks fail to measure, and models consequently fail to generate, truly robust and reliable code snippets. We introduce PACT, a program assessment and contract-adherence evaluation framework, to bridge this gap. PACT is the first framework designed to systematically evaluate and enhance contract-adherence in LLM-generated code snippets alongside functional correctness. PACT’s contributions are threefold: First, it provides a comprehensive test-suite corpus focused on contract violations, extending HumanEval+ and MBPP+. Second, it enables a systematic analysis of code generation under varied prompting conditions. This analysis demonstrates that augmenting prompts with contract-violating test cases significantly enhance a model’s ability to respect contracts compared to using contract description alone. Finally, it introduces novel metrics to rigorously quantify contract adherence in both test generation and code generation. By revealing critical errors that conventional benchmarks overlook, PACT provides the rigorous and interpretable metrics to evaluate the robustness of LLM-generated code snippets in both functionality and contract-adherence. Our code and data are available at https://github.com/suhanmen/PACT.
1 Introduction
Test case generation is particularly crucial since it validates correctness, reveals corner cases, and supports automated evaluation. Traditional approaches such as random testing (Cha et al., 2015), symbolic execution (King, 1976; Yoon & Cha, 2024; Cadar et al., 2008), and search-based software testing (Harman & McMinn, 2010; Formica et al., 2024; Burnim & Sen, 2008) have laid the groundwork. Then, various large language model (LLM)-driven methods have recently been investigated, including zero-shot prompting (Chen et al., 2021a; Jaremko et al., 2025), coverage-guided test generation (Ryan et al., 2024; Lemieux et al., 2023; Sapozhnikov et al., 2024), mutation-informed generation (Dakhel et al., 2024; Cajica et al., 2021), slicing-based decomposition (Wang et al., 2024), and satisfiability modulo theories (SMT) solver-based test generation (Peleska et al., 2011; Cadar et al., 2008).
Yet current test case generation approaches (Srivastava & Payer, 2021; Yoon & Cha, 2024) and the resulting benchmarks (Chen et al., 2021a) still judge the quality of both test cases and programs exclusively through pass@k (Chen et al., 2021a). In practice, real-world software is governed by contracts (Meyer, 1992)—the preconditions and input validation rules that restrict valid inputs and define expected behavior for ill-formed ones. As a black-box metric, pass@k measures correctness only via input-output correspondence on well-formed inputs, which are test cases that have already been filtered to comply with these contracts. Consequently, evaluating code snippets solely by well-formed input-output correspondence without considering contracts leads to an inaccurate assessment, as it overlooks whether the generated code snippets enforce pre-conditions and input validation checks explicitly stated or implicitly included in the specification (Liu et al., 2023).
Typically, most coding competitions such as International Collegiate Programming Contest  [^1] ICPC website: https://icpc.global/ (ICPC) or International Olympiad in Informatics  [^2] IOI website: https://ioinformatics.org/ provide specifications that state the constraints to consider when implementing the function for the task. This description includes not only the functional goal—such as computing a sum or sorting a list—but also a set of explicit contracts, such as input bounds, type restrictions, and required error-handling behaviors. These contracts often phrased in natural language or implied via examples, act as contracts that describe the valid input space and define the expected behavior under edge or invalid conditions. Figure 1 provides a clear illustration, showing how a simple functional task description contains hidden contracts that are often overlooked. For instance, a function designed to find duplicate values in a list of numbers, has_dup(nums), has an implicit contract that the input must be a list containing only integers. A correct implementation should therefore include specific assertions—such as assert isinstance(nums, list) and assert all(isinstance(x, int) for x in nums)—to verify the input’s type before proceeding with the functional logic. If such checks are omitted, the program may appear correct on well-formed inputs like [1, 2, 3]. However, it will fail to reject ill-formed inputs like ’a’ or [’2’, 2], thus failing to reflect the intended specification. Ignoring these contracts in code generation poses serious reliability and safety issues. For instance, a module might silently accept an invalid input, such as a negative number where a positive one is expected. By doing so, it can propagate a hidden fault by passing a nonsensical result to another part of the system. This chain reaction often leads to logical errors that are extremely difficult to trace back to their original source.
Figure 1: PACT’s contract-violating test uncovers an implicit constraint that conventional functional tests miss, proving the need for contract-aware evaluation.
Therefore, contracts are not optional safeguards but integral parts of the specification that define the boundary of valid behavior. Without them, even code snippets that appear functionally correct cannot be trusted to operate safely and reliably. Empirical studies reveal that LLM-generated code snippets frequently overlook these contracts and often fails to guard against ill-formed inputs that violate them. As a result, a generated code may incorrectly pass tests that it should fail, creating the illusion of correctness under default test cases. This shows that many seemingly correct solutions in fact ignore contractual requirements, underscoring that contract adherence is a fundamental necessity for trustworthy code generation.
We introduce PACT, a program assessment and contract-adherence evaluation framework, to bridge this gap. The main focus of PACT is twofold: first, to construct high-quality contract-violating test cases, and second, to use them for a systematic analysis of contract-aware code generation. For test case generation, our approach leverages an SMT solver to systematically explore diverse combinations of contract violations, moving beyond simplistic, single-violation tests.
For example, for a function with the contracts assert isinstance(r, float), assert r > 0, and assert h > 0, our method can generate an input such as r = -2.5 and h = 5. This test case is guaranteed to satisfy the type contract for r and the positivity contract for h, while precisely violating only the positivity contract for r. By pruning logically inconsistent combinations in advance, the solver ensures that only feasible and semantically valid test cases are generated. For our analysis of code generation, we then use these precisely targeted test cases to evaluate how different prompting strategies guide an LLM to produce implementations that properly enforce the intended contracts.
In summary, current benchmarks overlook contract violations, thereby inflating the perception of correctness. Addressing this limitation, we propose the PACT framework. PACT extends HumanEval+ and MBPP+ to generate test cases focused on contract violations and introduces novel metrics to quantify contract adherence.
2 Related work
2.1 Automated Test and Code Generation
Automated test and code generation have deep roots in traditional software engineering. Early test case generation was typically categorized into black-box methods like random and mutational fuzzing (Godefroid et al., 2005; Claessen & Hughes, 2000), white-box techniques such as symbolic execution (He et al., 2021; Cha et al., 2022; Godefroid et al., 2008), and grey-box approaches like coverage-guided fuzzing (Choi et al., 2019; Stephens et al., 2016; She et al., 2024; Qian et al., 2022). Similarly, early code generation relied on methods like probabilistic grammar-based frameworks (Bielik et al., 2016) and specialized language models (Feng et al., 2020). The advent of the Transformer architecture marked a paradigm shift, establishing LLMs as the dominant approach in both fields (Chen et al., 2021b). The development of modern Code LLMs now typically involves pre-training on code corpora, followed by instruction tuning (Ouyang et al., 2022) and refinement through more advanced techniques like reinforcement learning with execution feedback (Gehring et al., 2025). However, despite these advancements, the primary goal has remained the enhancement of functional correctness, typically measured by pass@k on benchmarks with well-formed inputs.
2.2 Contracts
The Design by Contract (DbC) paradigm (Meyer, 1992) argues that software reliability depends on explicitly stated preconditions, postconditions, and invariants, with run-time assertions as the enforcement mechanism. However, mainstream automated testing and existing test-generation benchmarks (Wang et al., 2025; Jain et al., 2025) often overlook latent defects from missing contract checks. While related work on failure-inducing test cases (Zhang et al., 2024; Zhong et al., 2025) is effective at causing general exceptions, this is not the same as verifying adherence to the precise semantics of a given contract. For instance, for a function that requires a list of positive numbers, a generic failure-inducing test might use None to cause a TypeError, but this does not verify the specific rule that all numbers must be positive. In contrast, a contract-aware test like [10, 20, -5] is designed to be caught precisely by an assertion checking for positive values. This distinction highlights the need for systematic methods that can precisely target formal contract specifications rather than just triggering arbitrary errors.
2.3 SMT Solver
SMT (Satisfiability Modulo Theories) solvers are powerful engines for determining the satisfiability of complex logical formulas across various theories, such as arithmetic and strings (Barrett & Tinelli, 2018). They typically interface using SMT-LIB, a standardized formal language, with Z3 being a widely adopted implementation in automated testing (de Moura & Bjørner, 2008). Traditionally, techniques like symbolic execution have employed SMT solvers to find inputs that satisfy path constraints, focusing on functional coverage over well-formed inputs. In this paradigm, constraints are used as admission checks to filter out invalid data rather than as explicit targets for evaluation. In contrast, our framework leverages SMT solvers to systematically generate ill-formed inputs that precisely violate formalized contracts, allowing us to rigorously test for robustness. A detailed example of how we formulate contracts is provided in Appendix D.
3 The Need for Contract-Aware Evaluation
3.1 Why Contracts Matter: Blind Spots in Current Benchmarks
Recent studies on code and test case generation (Korraprolu et al., 2025; Sung et al., 2025; Li et al., 2022) have dominantly relied on pass@k to evaluate functional correctness. While effective, these metrics, along with approaches that target failure-inducing inputs (Peng et al., 2018), share a common limitation: they primarily operate within the bounds of a program’s legal input space Dyck et al. (2023). This practice has led to a critical blind spot in popular benchmarks such as HumanEval+ and MBPP+. By design, these benchmarks explicitly filter out and discard any test case that violates a program’s pre-conditions (Austin et al., 2021). Consequently, the evaluation process certifies that a program behaves correctly on well-formed inputs, but it fails entirely to assess the program’s robustness against ill-formed ones. This results in an incomplete and often inflated assessment of code quality, praising solutions that may be superficially correct but are fundamentally fragile.
This is where the software contracts become essential. As explained in Section 1, contracts are the rules that define the boundary of valid behavior. They are not optional safeguards but a core component of trustworthy software that specifies how a program must identify and reject invalid data. By ignoring contract adherence, existing benchmarks overlook a crucial dimension of software reliability.
3.2 Contract-Based Test Paradigm
Merely enlarging a single pool of valid inputs cannot reveal whether a model understands the boundary of the specification it needs to satisfy. The main focus, therefore, lies in constructing contract-violating test cases, which systematically explore the extent to which models enforce contract rules. A contract-violating test case is an input that violates one or more predicates from a set of contracts while satisfying the remaining specification. The reference implementation is augmented with run-time assertions for every predicate in the contract set, and a candidate program passes such a test only when it raises an error consistent with this augmented reference. Introducing contract-violating inputs uncovers false negatives that purely functional tests overlook and provides a rigorous measure of whether a program properly enforces contractual rules.
3.3 Task Setup
Each benchmark task consists of a natural language description, a set of contracts, and a functional implementation. We use HumanEval+ and MBPP+, where contract predicates are provided as assertion literals outside the prompt and reference code. Our task concerns contract-violating test generation and contract-aware code generation. For test generation, we automatically construct a compact set of contract-violating test cases that target specific contract rules and remain feasible with respect to the remaining rules for each benchmark task. These tests are used for evaluating whether LLM-generated code appropriately follows contracts. For contract-aware code generation, we generate code under two prompt conditions, where the first condition is a contract specification (CS) prompt, which includes the functional description and a natural language paraphrase of the contracts. The second condition is an example-augmented specification (EAS) prompt, which is augmented with the contract-violating test cases.
Figure 2: Running example of PACT with an Example-Augmented Specification (EAS) prompt, which integrates Contract Specification (CS) and contract-violating test cases (CVTs) to enforce contract-aware code generation.
4 Methodology
We propose PACT, a program assessment and contract-adherence evaluation framework designed to systematically evaluate and enhance the ability of LLMs to generate contract-compliant code. This framework consists of two main stages. First, we generate contract-violating test cases to rigorously assess whether LLM-generated code snippets enforce both functional specifications and explicit contracts. Second is the systematic analysis of code generation, where we use these test cases under different prompting conditions to evaluate a model’s contract awareness in detail. Unlike prior approaches that rely solely on functionality-based evaluation, PACT extends the evaluation paradigm with contract-violation tests, enabling a more precise and reliable analysis of contract adherence.
4.1 Contract-Violating Test-Case Generation
Direct LLM-based generation is inadequate for contract-aware test case construction. It must consider all subsets of the contract set when producing violating inputs. If a task has $n$ contracts, the number of non-empty violation combinations is $2^{n}-1$. Direct prompting often misses required combinations and yields contradictory inputs that violate unintended contracts and fails to ensure feasibility under the specification. LLMs also lack a built-in mechanism to verify that exactly the targeted contracts are violated while all others hold. We design an SMT-based approach to make these checks accurate and efficient. Our generation procedure is a two-step pipeline for contract-aware test case generation. First, an inference model translates natural language contracts into an Algebraic Data Type (ADT) program. This ADT format is the critical component: it provides a rigid, formal schema for the complex and often nested constraints found in contracts. This ensures that the subsequently generated rules for an SMT solver are syntactically valid and semantically precise. Second, given the ADT, the SMT solver constructs contract violation test cases (CVTs) by breaking one or more specified contracts while ensuring that all remaining contracts are satisfied. The solver validates satisfiability for each candidate combination and extracts concrete models—specific assignments of values that satisfy the constraints— to instantiate inputs for every valid case. We provide a running example of this procedure in Appendix D.
4.2 Contract Metrics for Test Case
Unlike standard test cases for functional correctness, test cases for contracts should be the negative samples, violating the contracts and triggering the corresponding assertions. As the concept of contract-violating test cases is different from the standard, we design metrics to analyze whether the generated test cases appropriately correspond to input contracts.
Let $A=\{a_{1},\ldots,a_{n}\}$ be the set of all contract assertions and $T=\{t_{1},\ldots,t_{m}\}$ be the generated test cases. $F_{t}\subseteq A$ is the set of violated assertions when executing a test case $t\in T$. Finally, $T_{neg}\subseteq T$ represents the set of test cases that successfully violated at least one contract.
Assert Violation Coverage (AVC)
AVC quantifies the coverage of assert statements that are successfully violated by all the test cases.
$AVC=\frac{|\{a_{i}\mid\exists t\in T_{neg}:a_{i}\in F_{t}\}|}{n}.$
The value of 1.0 for AVC ensures that all contract assertions are captured by the test cases. Lower values expose the unexplored regions of the contracts from the test cases.
Target Specificity (TS)
TS evaluates the precision of each test case by measuring how accurately it violates its intended set of target contracts. To formalize this, we first define $V\subseteq A$ as the set of contracts that a given test case $t\in T_{neg}$ is intended to violate. We then measure the alignment between this intended set $V_{t}$ and the actually violated set $F_{t}$ using the Jaccard Index. The final TS score is the average of these individual precision scores across all negative test cases:
$TS=\frac{1}{|T_{neg}|}\sum_{t\in T_{neg}}\frac{|F_{t}\cap V_{t}|}{|F_{t}\cup V_{t}|}.$
A score of 1.0 indicates perfect precision, where every test case violates exactly the set of contracts it was designed to. Lower scores reveal a discrepancy, indicating that test cases either failed to trigger their intended violations or caused unintended, collateral violations.
4.3 Contract-Aware Code Generation
This section details our methodology for systematically evaluating and enhancing the ability of LLMs to generate code snippets that robustly enforce contracts. To achieve a fine-grained understanding of a model’s contract awareness, we investigate its performance under two distinct prompting strategies. First, we establish a baseline condition, Contract Specification (CS). For this strategy, we use a powerful Commercial LLM to naturally integrate the contract rules from the HumanEval+ and MBPP+ datasets into the main text of the prompt. This creates a comprehensive prompt that describes both the functional goal and its contractual constraints in natural language. Second, we introduce an enhanced condition, which we term Example-Augmented Specification (EAS). This strategy builds upon the CS prompt by augmenting it with a single, precisely targeted contract- violating test cases for each described contract. This provides a concrete example of what constitutes a violation, intended to guide the model toward more robust code generation. A concrete example of this EAS prompt is shown in Figure 2.
To measure the impact of these prompting strategies, we assess the generated code snippets using a comprehensive suite of metrics for both functional correctness and contract adherence. Functional correctness is measured using the standard pass@k metric over a set of valid test cases. Contract adherence is evaluated from two perspectives: runtime enforcement using test cases and static alignment of the generated code snippets with the ground-truth contracts, as detailed in Section 4.4.
4.4 Contract Metric for Code Generation
For the evaluation of the generated code snippets, we employ pass@k to measure functional correctness on valid inputs and use AVC to measure the correctness of the generated contract assertions. We also design two additional metrics to more precisely evaluate contract-aware code generation.
Let $A=\{a_{1},\ldots,a_{n}\}$ be the set of ground-truth contract assertions, and let $\hat{A}=\{\hat{a}_{1},\ldots,\hat{a}_{m}\}$ be the set of assertions extracted from the LLM-generated code. Let $M\subseteq A\times\hat{A}$ be the set of pairs $(a_{i},\hat{a}_{j})$ where the ground-truth contract $a_{i}$ and the generated assertion $\hat{a}_{j}$ are determined to be semantically equivalent.
Assertion Alignment Recall (AAR)
AAR measures the model’s ability to implement all required contracts without omission. It is the proportion of ground-truth contracts that are successfully covered by at least one assertion in the generated code, functioning as a recall metric.
$AAR=\frac{|\{a_{i}\in A\mid\exists\hat{a}_{j}\in\hat{A}:(a_{i},\hat{a}_{j})\in M\}|}{n}.$
A high AAR score indicates that an output code ensures that all required contract specifications are generated.
Assertion Alignment Precision (AAP)
AAP measures the accuracy of the generated assertions, penalizing irrelevant or hallucinated ones. It is the proportion of generated assertions that correspond to a valid ground-truth contract, functioning as a precision metric.
$AAP=\frac{|\{\hat{a}_{j}\in\hat{A}\mid\exists a_{i}\in A:(a_{i},\hat{a}_{j})\in M\}|}{m}.$
A high AAP score indicates that the code does not contain assertions of unnecessary or incorrect checks.
5 Experimental settings
5.1 Datasets
Our study utilizes HumanEval+ and MBPP+ benchmarks (Liu et al., 2023), which are enriched with assertion-level contracts. However, these benchmarks have a critical limitation for evaluating code robustness: the contract specifications are neither included in the model prompts, nor are contract-violating inputs included in the official test suites. This evaluation setup exclusively tests for functional correctness on well-formed inputs, leading to an inflated perception of code quality. We construct a supplementary dataset containing contract-violating test cases, as described in Section 4.1.
5.2 Candidate Models
We utilize o4-mini as our primary test case generator. For our code generation experiments, we evaluate a set of open-source models including gemma-3-12B-it (gemma-3), Deepseek-R1-Distill-Qwen-14B (DeepSeek-R1), Qwen3-14B (Qwen-3), and Phi-4-reasoning-plus (Phi-4-plus).
6 Empirical Studies
Our empirical study is structured over three research questions (RQs) designed to evaluate the PACT framework from multiple perspectives. The evaluation is conducted on HumanEval+ and MBPP+. We begin by assessing the quality and precision of the CVTs generated by PACT (RQ1). Next, we investigate whether providing these concrete test cases is more effective for eliciting contract-awareness than using abstract natural language descriptions alone (RQ2). Finally, we analyze the resulting trade-off between the enforced contract adherence and the functional correctness of the generated code snippets (RQ3).
Table 1: Evaluation results of the test case generation on HumanEval+ and MBPP+.
Benchmark Method AVC ($\uparrow$}) TS ($\uparrow$}) AVG ($\uparrow$})
HumanEval+ o4-mini 97.14% 75.60% 86.37%
o4-mini + SMT Solver 95.53% 85.81% 90.67%
MBPP+ o4-mini 94.67% 69.11% 81.89%
o4-mini + SMT Solver 93.50% 84.54% 89.02%
6.1 RQ1: How Effective is PACT in Generating High-Quality Contract- Violating Test Cases?
Our first research question investigates the effectiveness of our proposed framework, PACT, in generating high-quality CVTs. A naive baseline approach, which we refer to as o4-mini in Table 1, uses an LLM o4-mini for direct test case generation from given contracts. This method is often inadequate, as it disregards dependencies among contracts and produces logically inconsistent violations, making it difficult to precisely verify individual contract predicates.
In contrast, PACT (o4-mini+ SMT Solver) employs a more robust two-stage approach. First, an LLM converts contracts into rules for an SMT solver. Subsequently, an SMT solver uses these rules to generate test cases. This process guarantees the generation of logically sound test cases that are precisely targeted to violate a specific subset of contracts while adhering to the rest.
The empirical results in Table 1 validate the superiority of PACT. While both methods achieve high AVC scores, PACT significantly outperforms the direct generation in TS, achieving over 10%p increases on both HumanEval+ and MBPP+. This demonstrates that our two-stage approach successfully gives test cases precisely for intended contracts. The baseline’s slightly higher AVC is an expected outcome, as direct generation often triggers a wide range of assertion errors indiscriminately, artificially inflating coverage. PACT, on the other hand, focuses on valid, targeted violations.
While PACT generates valid rules, providing a precise test for the input contracts, there are still errors. A case-specific analysis of minor errors that occur even with well-formed rules is detailed in Appendix A.
Table 2: Evaluation results of contract adherence in code generation on HumanEval+.
Model Mode pass@1 ($\uparrow$}) AVC ($\uparrow$}) AAR ($\uparrow$}) AAP ($\uparrow$}) AVG ($\uparrow$})
gemma-3 CS 84.41% 24.85% 11.41% 14.04% 32.79%
EAS 82.94% 91.02% 28.07% 27.77% 57.45%
DeepSeek CS 73.78% 44.12% 15.65% 16.97% 37.63%
EAS 71.77% 79.29% 27.62% 28.01% 51.67%
Qwen3 CS 78.92% 28.04% 13.17% 22.55% 35.67%
EAS 77.83% 87.81% 31.53% 36.09% 58.31%
Phi-4-plus CS 72.23% 52.91% 18.78% 21.09% 41.25%
EAS 67.08% 69.50% 21.33% 20.06% 44.49%
Table 3: Evaluation results of contract adherence in code generation on MBPP+.
Model Mode pass@1 ($\uparrow$}) AVC ($\uparrow$}) AAR ($\uparrow$}) AAP ($\uparrow$}) AVG ($\uparrow$})
gemma-3 CS 78.56% 57.99% 17.50% 17.93% 41.49%
EAS 78.60% 95.57% 32.29% 31.82% 59.57%
DeepSeek CS 62.53% 64.20% 17.59% 17.57% 40.47%
EAS 60.15% 86.70% 28.23% 27.94% 45.47%
Qwen3 CS 72.41% 70.86% 21.09% 22.99% 46.84%
EAS 72.63% 94.85% 31.54% 32.30% 57.83%
Phi-4-plus CS 64.89% 67.33% 24.26% 24.65% 45.28%
EAS 63.76% 74.88% 29.20% 28.95% 49.20%
6.2 RQ2: Are Concrete Test Cases More Effective than Abstract Descriptions for Eliciting Contract-Awareness?
We compared two prompting methods across various models in Tables 2 and 3. CS, our baseline, provides only an abstract natural language description of the contracts. In contrast, EAS augments the CVTs, generated from Section 4.1, to the CS prompt.
The results demonstrate that for all of our base models, EAS is a more powerful method for eliciting contract-aware code generation. EAS achieves 33.7%p, 11.3%p, and 9.4%p improvements for AVC, AAR, and AAP, respectively, over CS. For models such as Qwen3, DeepSeek-R1, and Phi-4-plus, switching from CS to EAS leads to a dramatic improvement across all contract adherence metrics. This indicates that concrete examples of CVTs provide a clear and unambiguous signal that forces the model to move beyond purely functional logic. The increase in AAR and AAP indicates that the models are correctly generating test cases required for assertion checks, while the relatively huge improvements in AVC show that the resulting codes are more robust against a wider range of invalid inputs.
Concrete examples of contract violations provide a clear, unambiguous signal that forces the model to move beyond purely functional logic and implement the explicit enforcement of contracts—the predefined agreements on how ill-formed inputs must be rejected. For concrete examples of the code snippets and their corresponding LLM-generated test cases, please refer to Appendix B. However, this heightened focus on contract enforcement introduces a notable trade-off, as the model may generate more complex code that impacts its performance on purely functional correctness metrics. This trade-off is analyzed in further detail in the following section RQ3.
6.3 RQ3: How Does Enforcing Contract Adherence Impact the Functional Correctness of Generated Code?
While the RQ2 established that augmenting prompts with CVTs significantly enhances a model’s contract adherence, this section investigates the resulting trade-off with respect to its functional correctness. Our analysis reveals that a heightened focus on contract enforcement often comes at the cost of a measurable decline in performance on standard functional correctness benchmarks. This trade-off can be attributed to a shift in the model’s optimization focus. When provided with only a functional description, the LLM’s sole objective is to produce code snippets that pass functionality-focused tests. However, when the prompt is augmented with CVTs, the model must simultaneously satisfy two competing objectives: generating functionally correct logic and enforcing the specified contractual preconditions. This dual objective increases the complexity of the code generation task. The model is compelled to allocate part of its reasoning capacity to interpreting and implementing the contract rules, which can lead to subtle errors or oversights in the core functional logic. A concrete example of this is presented in Appendix C. Consequently, while the resulting code snippets are more robust and secure against invalid inputs, they may exhibit a lower pass@1 rate on test suites composed entirely of valid, well-formed inputs.
In conclusion, our findings reveal an inherent tension between contract adherence and functional correctness in LLM-based code generation. The pursuit of robustness through contract enforcement comes at a tangible cost to functionality. This trade-off highlights the critical need for advanced training paradigms—such as the reinforcement learning with multi-objective rewards proposed by PACT—capable of optimizing for both objectives simultaneously.
7 Conclusion
We introduce PACT, the first framework to redefine code and test case correctness by evaluating adherence to task specifications through both functionality and contract-based behavior. While prior benchmarks assess only pass@k on well-formed inputs, PACT introduces a comprehensive paradigm with dual test suites—one for functionality and one for contract violations—along with specific metrics to analyze contract awareness and uncover latent defects, enabling a more principled evaluation of code robustness. Our empirical evaluation demonstrates the effectiveness of PACT across multiple dimensions. We first show that PACT’s SMT-solver-based test case generation method ensures more accurate CVTs than direct generation with over 10%p better performance. Furthermore, our results reveal that augmenting prompts with these CVTs is a highly effective strategy for generating robust and contract-aware code, achieving 18.13%p increase in contract-specific metrics and 12.8%p in total average. Our empirical studies on various LLMs demonstrate the effectiveness of PACT, achieving approximately, but also uncovers a critical trade-off between this enhanced contract adherence and a model’s performance on functional correctness. These findings confirm that PACT provides a more complete and realistic assessment of an LLM’s contract-aware code generation capabilities, moving beyond the limitations of existing benchmarks.
8 Future work directions
While this work focuses on evaluating contract adherence, a natural next step is to actively improve it through advanced training methodologies. A promising direction is to leverage the novel metrics introduced in our PACT framework as direct training signals. The novel metrics introduced in our PACT framework, such as the runtime metric Assert Violation Coverage (AVC) and the static metrics Assertion Alignment Recall (AAR) and Precision (AAP), are particularly well-suited for this purpose and could be integrated into a multi-objective reward function for RL. This approach could enable models to learn to navigate the trade-off between functional correctness and contract adherence more effectively, optimizing for both objectives simultaneously.
Appendix A Case Study: Logical Contradictions in Direct LLM Test Case Generation
HumanEval
In Figure 3 shown in the code snippet, this task includes three sequential contracts: assert_0 checks if the input is a list, assert_1 verifies that all elements in the list are strings, and assert_2 ensures that all strings consist only of digits. A critical dependency exists between these contracts. Specifically, assert_2 can only be evaluated if assert_1 is satisfied, because the isdigit() method is only valid for string types. A test case designed to violate assert_1 while satisfying assert_0 and assert_2 would therefore be a logically contradictory combination, as a non-string element would cause a TypeError before assert_2 could be checked. Despite this, a direct LLM generation approach often produces such invalid combinations. For instance, when tasked to generate test cases, the LLM produces inputs such as [123, "456"], ["789", [0]], and ["456", false]. These examples fail to isolate a specific contract violation. This highlights a fundamental weakness of the approach, as the LLM tends to generate simplistic contract-violation test cases that fail to respect the logical relationships among contracts.
Mbpp
In Figure 4 shown in the code snippet, this task includes four main contracts, which can be grouped by their dependency. The initial contracts, assert_0 and assert_1, perform type checking to verify that both inputs are of a numeric type, such as an integer or a floating-point number. The subsequent contracts, assert_2 and assert_3, check numeric properties, such as ensuring the numbers are positive or fall within a specific range. A critical dependency exists between these groups of contracts. Specifically, the numeric property checks in assert_2 and assert_3 can only be evaluated if the type checks in assert_0 and assert_1 are satisfied. For example, a non-numeric type like a string or null cannot be evaluated for properties like being positive. Therefore, creating a contract-violation test case that violates the initial type contracts (assert_0 or assert_1) while simultaneously satisfying the subsequent property contracts (assert_2 and assert_3) is a logical impossibility. Despite this, a direct LLM generation approach often produces such logically flawed combinations. For instance, when tasked to generate test cases, the LLM produces inputs such as ["abc", null], [null, "abc"], [[1], {"x":1}], and [{"r":1}, [2]]. Crucially, while these examples successfully violate the initial type contracts, they all inherently fail to satisfy assert_2 and assert_3, yet they are generated as if such a combination were possible. This highlights a fundamental weakness of the approach, as the LLM tends to generate simplistic contract-violation test cases that fail to respect the logical relationships among contracts.
HumanEval/113 def odd_count(lst): assert type(lst) == list, "invalid inputs" # \$_CONTRACT_\$ assert all(isinstance(s, str) for s in lst), "invalid inputs" # \$_CONTRACT_\$ assert all(s.isdigit() for s in lst), "invalid inputs" # \$_CONTRACT_\$ ans, template = [], "the number of odd elements in the string i of the input." for s in lst: odd_cnt = len(list(filter(lambda ch: int(ch) % 2 == 1, s))) ans.append(template.replace("i", str(odd_cnt))) return ans """ Contract List: assert_0: assert type(lst) == list, "invalid inputs assert_1: assert all(isinstance(s, str) for s in lst), "invalid inputs assert_2: assert all(s.isdigit() for s in lst), "invalid inputs """"
Figure 3: Code and contracts for HumanEval.
MBPP/731 import math def lateralsurface_cone(r,h): assert isinstance(r, (int, float)), "invalid inputs" # \$_CONTRACT_\$ assert isinstance(h, (int, float)), "invalid inputs" # \$_CONTRACT_\$ assert r > 0, "invalid inputs" # \$_CONTRACT_\$ assert h > 0, "invalid inputs" # \$_CONTRACT_\$ l = math.sqrt(r * r + h * h) return math.pi * r * l """ Contract List: assert_0: assert isinstance(r, (int, float)), "invalid inputs assert_1: assert isinstance(h, (int, float)), "invalid inputs assert_2: assert r > 0, "invalid inputs assert_3: assert h > 0, "invalid inputs """
Figure 4: Code and contracts for MBPP.
Appendix B Case Study: Contract Enforcement under Different Prompting Conditions
We present a direct comparison of code snippets that the DeepSeek model generated for the MBPP/11 task under two prompting conditions. This comparison illustrates why providing CVTCs is more effective than relying on natural language descriptions alone.
Figure 5 shows the code snippets produced when the model received only the natural language prompt. The prompt contains enough information to infer all necessary contracts. It specifies two explicit contracts, namely that the first input must be a non empty string and that the second input must be a string of length one.It also implies two type contracts that both inputs must be strings. The generated code snippet correctly implements the explicit length based contracts with assert len(s) > 0 and assert len(char) == 1. However, it omits the implicit type checks and treats the word “string” as descriptive context rather than a strict precondition. As a result, the function fails to enforce the contracts under non string inputs.
In contrast, Figure 6 displays the output when the prompt was augmented with concrete CVTCs such as remove_Occ(2.0, 3.0). These test cases provide an unambiguous and executable signal that non string inputs are invalid and must be rejected. The model generates a more robust implementation that now includes the crucial isinstance(s, str) and isinstance(char, str) assertions.
This comparison shows that natural language descriptions may contain all required information, yet their inherent ambiguity can prevent the model from adhering to the contract. CVTCs make these implicit contracts explicit and enforceable. They compel the model to produce code snippet that is more comprehensive and adheres to the intended contracts.
MBPP/11 """ Write a python function that removes the first and last occurrence of a given character from the string; the first input must be a non-empty string and the second input must be a string of length one. """ def remove_Occ(s, char): assert len(s) > 0, "First input must be a non-empty string." assert len(char) == 1, "Second input must be a string of length one." first = s.find(char) last = s.rfind(char) if first == -1: return s return s[:first] + s[first+1:last] + s[last+1:]
Figure 5: Code generated by DeepSeek with the contract specification (CS) prompt.
MBPP/11 """ Write a python function that removes the first and last occurrence of a given character from the string; the first input must be a non-empty string and the second input must be a string of length one. # Contract Test Cases: >>> remove_Occ(2.0, 3.0) "AssertionError: invalid input" >>> remove_Occ(5.0, 4.0) "AssertionError: invalid input" >>> remove_Occ(’’, 2.0) "AssertionError: invalid input" >>> remove_Occ(2.0, ’’) "AssertionError: invalid input" """ def remove_Occ(s, char): assert isinstance(s, str) and len(s) > 0, "invalid input" assert isinstance(char, str) and len(char) == 1, "invalid input" first_occ = s.find(char) last_occ = s.rfind(char) if first_occ == -1: return s return s[:first_occ] + s[first_occ+1:last_occ] + s[last_occ+1:]
Figure 6: Code generated by DeepSeek with the example-augmented specification (EAS) prompt.
Appendix C Case Study: Contract Adherence vs. Functional Correctness
A case study from the HumanEval/142 task provides a concrete example of the trade-off between contract adherence and functional correctness. The figures 79 compare the ground-truth implementation (golden label), the code generated from a contract natural language prompt (CS), and the code generated from a prompt augmented with contract-violating test cases (CVTs).
Figure 8 displays the code generated from the NL-only prompt, which describes both functional requirements and contractual constraints. It fails to translate the natural language contract descriptions into the necessary assertions, it successfully generates code snippets that passes all standard functionality tests. This outcome suggests the model prioritized achieving functional correctness—even developing a logic for handling multiples of 12—at the complete expense of implementing the required robustness checks, leaving the code snippets vulnerable to invalid inputs.
In contrast, Figure 9 shows the code generated when the prompt was augmented with CVTCs. Here, the model successfully implements the necessary assertions, demonstrating strong contract adherence. However, this heightened focus on contract enforcement directly compromises its functional correctness. The model generates a flawed conditional logic that explicitly excludes common multiples of 3 and 4. For an index like 0 or 12, the code incorrectly defaults to the else block, failing to square the number as required by the specification in the Golden Label (Figure 7).
This example clearly illustrates the trade-off. When compelled to handle the dual objectives of implementing contracts and functional logic, the model’s reasoning capacity is divided. The added complexity of enforcing contracts leads to a critical oversight in the core functional logic, resulting in code snippet that is robust but functionally incorrect. This underscores the inherent tension between these two competing objectives in LLM-based code generation.
HumanEval/142 def sum_squares(lst): assert type(lst) == list, "invalid inputs" # \$_CONTRACT_\$ assert all(type(x) == int for x in lst), "invalid inputs" # \$_CONTRACT_\$ ans = 0 for i, num in enumerate(lst): if i % 3 == 0: ans += num ** 2 elif i % 4 == 0: ans += num ** 3 else: ans += num return ans
Figure 7: The ground-truth implementation, correct in both functionality and contracts.
HumanEval/142 def sum_squares(lst): total = 0 for i in range(len(lst)): num = lst[i] if i % 3 == 0 and i % 4 == 0: total += num ** 2 elif i % 3 == 0: total += num ** 2 elif i % 4 == 0: total += num ** 3 else: total += num return total
Figure 8: A functionally correct implementation that fails to enforce contracts.
HumanEval/142 def sum_squares(lst): assert isinstance(lst, list), "AssertionError: invalid input" for elem in lst: assert isinstance(elem, int), "AssertionError: invalid input" total = 0 for index, num in enumerate(lst): if index % 3 == 0 and index % 4 != 0: total += num ** 2 elif index % 4 == 0 and index % 3 != 0: total += num ** 3 else: total += num return total
Figure 9: A robust implementation that enforces contracts but fails on functionality.
Appendix D Formalizing Contracts into SMT-LIB: Template and a Case Study
This section details the structure of the SMT-LIB template used by our PACT framework. SMT-LIB is a standardized, text-based language used to interface with SMT solvers. It provides a formal syntax for declaring variables, defining functions, and asserting logical formulas, allowing complex problems to be translated into a format that a solver can systematically analyze for satisfiability. Our framework leverages this language to translate nuanced, natural language contracts into a formal representation that can be reasoned about with logical precision.
Figure 10 shows the base template we designed for this purpose. It is composed of several key components, each serving a distinct role in the test generation process. The placeholders within this template are populated as follows:
  • CANONICAL PYTHON-LIKE ADT: This fixed block defines a universal data structure for representing common Python types. This allows the SMT solver to reason about various input types in a standardized way.
  • HELPER FUNCTIONS: This section is populated with custom functions needed to define the contracts for a specific task. For example, a function to check if a string contains only digits would be defined here.
  • INPUT: The input variables for the function under test are declared here.
  • BASIC_STRUCTURE: This section defines fundamental structural constraints on the inputs, such as ensuring a variable is a list composed of integer values.
  • CONTRACT_DEFS: The specific logical rules of each contract are translated into formal predicates in this section.
  • COMBINATION: This is the core logic for generating a test case. It contains assertions stating which contracts must be satisfied and which must be violated. The SMT solver then attempts to find a concrete model that satisfies this exact combination of constraints.
Figure 11 shows the ground-truth Python implementation for the HumanEval/11 task, which requires a function that takes two binary strings of equal length. The SMT-LIB formalization of these requirements is shown in Figure 12. The three assert statements in the Python code directly correspond to the three formal contracts defined in SMT-LIB:
  • C0 verifies that both inputs are strings, corresponding to the assertion assert isinstance(a, str) and isinstance(b, str).
  • C1 ensures their lengths are equal, corresponding to assert len(a) == len(b).
  • C2 checks that they are valid binary strings using a custom isBinaryString helper function, corresponding to assert set(a).issubset("0", "1") and set(b).issubset("0", "1").
The COMBINATION block determines the goal of the test case generation. By choosing to assert either the contract itself, such as (assert (C0)), or its negation, such as (assert (not C0)), for each rule, this block can instruct the SMT solver to find a test case for any desired combination of contract satisfactions and violations. The specific instance in the figure 12, for example, asserts the negation of all three contracts to generate a test case that violates every precondition simultaneously.
The SMT-LIB template ADT_BASE_TEMPLATE = """ (set-logic ALL) ; ==== CANONICAL PYTHON-LIKE ADT (DO NOT MODIFY) ==== (declare-datatypes ((Value 0)) ( ((IntVal (ival Int)) (FloatVal (fval Real)) (StrVal (sval String)) (BoolVal (bval Bool)) (Nil) (Cons (head Value) (tail Value))) )) ; === ADD HELPER FUNCTIONS HERE === <<HELPER_FUNCTIONS>> ; === Inputs === <<INPUT>> ; === BASIC STRUCTURE === <<BASIC_STRUCTURE>> ; === Contract predicates === <<CONTRACT_DEFS>> ; === COMBINATION === <<COMBINATION>> (check-sat) (get-model) """
Figure 10: The SMT-LIB template used for formalizing contracts.
HumanEval/11 from typing import List def string_xor(a: str, b: str) -> str: assert isinstance(a, str) and isinstance(b, str), "invalid inputs" # \$_CONTRACT_\$ assert len(a) == len(b), "invalid inputs" # \$_CONTRACT_\$ assert set(a).issubset({"0", "1"}) and set(b).issubset({"0", "1"}), "invalid inputs" # \$_CONTRACT_\$ return "".join(str(int(a[i]) ^ int(b[i])) for i in range(len(a)))
Figure 11: The ground-truth implementation in HumanEval/11
The SMT-LIB template ADT_BASE_TEMPLATE = """ (set-logic ALL) ; ==== CANONICAL PYTHON-LIKE ADT (DO NOT MODIFY) ==== (declare-datatypes ((Value 0)) ( ((IntVal (ival Int)) (FloatVal (fval Real)) (StrVal (sval String)) (BoolVal (bval Bool)) (Nil) (Cons (head Value) (tail Value))) )) ; === ADD HELPER FUNCTIONS HERE === (define-fun Safe_Sval ((x Value)) String (ite (is-StrVal x) (sval x) "")) (define-fun isBinaryString $($s Value)) Bool (and (is-StrVal s) (str.in.re (Safe_Sval s) (re.* (re.union (str.to.re "0") (str.to.re "1")))))) ; === Inputs === (declare-const a Value) (declare-const b Value) ; === BASIC STRUCTURE === ; === Contract predicates === (define-fun C0 () Bool (and (is-StrVal a) (is-StrVal b))) (define-fun C1 () Bool (= (str.len (Safe_Sval a)) (str.len (Safe_Sval b)))) (define-fun C2 () Bool (and (isBinaryString a) (isBinaryString b))) ; === COMBINATION === (assert (not C0) (assert (not C1) (assert (not C2) (check-sat) (get-model) """
Figure 12: An example of the SMT-LIB template populated for HumanEval/11

### v3

@@
+ContractEval: A Benchmark for Evaluating Contract-Satisfying Assertions in Code Generation
+Abstract
+Current code generation benchmarks measure functional correctness on well-formed inputs, as test cases are curated to satisfy input preconditions.
+This leaves a gap: generated programs may appear correct but fail to satisfy contracts—assertion-level validity constraints for rejecting ill-formed inputs.
+We introduce ContractEval, a benchmark for evaluating contract-satisfying assertions in code generation, i.e., whether code rejects contract-violating inputs by triggering intended assertions.
+Built on HumanEval+ and MBPP+, ContractEval augments each task with contract-violation tests derived from reference assertions.
+We synthesize these via a neuro-symbolic pipeline: an LLM converts assertion clauses into constraints, and an SMT solver enumerates satisfiable violation combinations to generate inputs that violate selected clauses while satisfying the rest.
+Across five code LLMs, standard prompting yields 0% contract satisfaction on contract-violating test cases.
+Adding a small number of contract-violation examples increases contract satisfaction to 49–53% and retains 92% of the original pass@1.
+Our code is available at https://github.com/suhanmen/ContractEval.
+[…]
 envname-P envname#1
-Abstract
-Prevailing code generation benchmarks, such as HumanEval+ and MBPP+, primarily evaluate large language models (LLMs) with pass@k on functional correctness using well-formed inputs.
-However, they ignore a crucial aspect of real-world software: adherence to contracts—the preconditions and validity constraints that dictate how ill-formed inputs must be rejected.
-This critical oversight means that existing benchmarks fail to measure, and models consequently fail to generate, truly robust and reliable code snippets.
-We introduce PACT, a program assessment and contract-adherence evaluation framework, to bridge this gap.
-PACT is the first framework designed to systematically evaluate and enhance contract-adherence in LLM-generated code snippets alongside functional correctness.
-PACT’s contributions are threefold: First, it provides a comprehensive test-suite corpus focused on contract violations, extending HumanEval+ and MBPP+.
-Second, it enables a systematic analysis of code generation under varied prompting conditions.
-This analysis demonstrates that augmenting prompts with contract-violating test cases significantly enhance a model’s ability to respect contracts compared to using contract description alone.
-Finally, it introduces novel metrics to rigorously quantify contract adherence in both test generation and code generation.
-By revealing critical errors that conventional benchmarks overlook, PACT provides the rigorous and interpretable metrics to evaluate the robustness of LLM-generated code snippets in both functionality and contract-adherence.
-Our code and data are available at https://github.com/suhanmen/PACT.
+ContractEval: A Benchmark for Evaluating Contract-Satisfying Assertions in Code Generation
+1Soohan Limthanks: Equal contribution.
+,   1Joonghyuk Hahn11footnotemark: 1  ,   2Hyunwoo Park  ,   2Sang-Ki Ko  ,   1Yo-Sub Hanthanks: Corresponding author.
+1Yonsei University, Seoul, Republic of Korea  2University of Seoul, Seoul, Republic of Korea {aness1219,greghahn,emmous}@yonsei.ac.kr , {hwpark03,sangkiko}@uos.ac.kr
 1 Introduction
-Test case generation is particularly crucial since it validates correctness, reveals corner cases, and supports automated evaluation.
-Traditional approaches such as random testing (Cha et al., 2015), symbolic execution (King, 1976; Yoon & Cha, 2024; Cadar et al., 2008), and search-based software testing (Harman & McMinn, 2010; Formica et al., 2024; Burnim & Sen, 2008) have laid the groundwork.
-Then, various large language model (LLM)-driven methods have recently been investigated, including zero-shot prompting (Chen et al., 2021a; Jaremko et al., 2025), coverage-guided test generation (Ryan et al., 2024; Lemieux et al., 2023; Sapozhnikov et al., 2024), mutation-informed generation (Dakhel et al., 2024; Cajica et al., 2021), slicing-based decomposition (Wang et al., 2024), and satisfiability modulo theories (SMT) solver-based test generation (Peleska et al., 2011; Cadar et al., 2008).
-Yet current test case generation approaches (Srivastava & Payer, 2021; Yoon & Cha, 2024) and the resulting benchmarks (Chen et al., 2021a) still judge the quality of both test cases and programs exclusively through pass@k (Chen et al., 2021a).
-In practice, real-world software is governed by contracts (Meyer, 1992)—the preconditions and input validation rules that restrict valid inputs and define expected behavior for ill-formed ones.
-As a black-box metric, pass@k measures correctness only via input-output correspondence on well-formed inputs, which are test cases that have already been filtered to comply with these contracts.
-Consequently, evaluating code snippets solely by well-formed input-output correspondence without considering contracts leads to an inaccurate assessment, as it overlooks whether the generated code snippets enforce pre-conditions and input validation checks explicitly stated or implicitly included in the specification (Liu et al., 2023).
-Typically, most coding competitions such as International Collegiate Programming Contest  [^1] ICPC website: https://icpc.global/ (ICPC) or International Olympiad in Informatics  [^2] IOI website: https://ioinformatics.org/ provide specifications that state the constraints to consider when implementing the function for the task.
-This description includes not only the functional goal—such as computing a sum or sorting a list—but also a set of explicit contracts, such as input bounds, type restrictions, and required error-handling behaviors.
-These contracts often phrased in natural language or implied via examples, act as contracts that describe the valid input space and define the expected behavior under edge or invalid conditions.
-Figure 1 provides a clear illustration, showing how a simple functional task description contains hidden contracts that are often overlooked.
-For instance, a function designed to find duplicate values in a list of numbers, has_dup(nums), has an implicit contract that the input must be a list containing only integers.
-A correct implementation should therefore include specific assertions—such as assert isinstance(nums, list) and assert all(isinstance(x, int) for x in nums)—to verify the input’s type before proceeding with the functional logic.
-If such checks are omitted, the program may appear correct on well-formed inputs like [1, 2, 3].
-However, it will fail to reject ill-formed inputs like ’a’ or [’2’, 2], thus failing to reflect the intended specification.
-Ignoring these contracts in code generation poses serious reliability and safety issues.
-For instance, a module might silently accept an invalid input, such as a negative number where a positive one is expected.
-By doing so, it can propagate a hidden fault by passing a nonsensical result to another part of the system.
-This chain reaction often leads to logical errors that are extremely difficult to trace back to their original source.
-Figure 1: PACT’s contract-violating test uncovers an implicit constraint that conventional functional tests miss, proving the need for contract-aware evaluation.
-Therefore, contracts are not optional safeguards but integral parts of the specification that define the boundary of valid behavior.
-Without them, even code snippets that appear functionally correct cannot be trusted to operate safely and reliably.
-Empirical studies reveal that LLM-generated code snippets frequently overlook these contracts and often fails to guard against ill-formed inputs that violate them.
-As a result, a generated code may incorrectly pass tests that it should fail, creating the illusion of correctness under default test cases.
-This shows that many seemingly correct solutions in fact ignore contractual requirements, underscoring that contract adherence is a fundamental necessity for trustworthy code generation.
-We introduce PACT, a program assessment and contract-adherence evaluation framework, to bridge this gap.
-The main focus of PACT is twofold: first, to construct high-quality contract-violating test cases, and second, to use them for a systematic analysis of contract-aware code generation.
-For test case generation, our approach leverages an SMT solver to systematically explore diverse combinations of contract violations, moving beyond simplistic, single-violation tests.
-For example, for a function with the contracts assert isinstance(r, float), assert r > 0, and assert h > 0, our method can generate an input such as r = -2.5 and h = 5.
-This test case is guaranteed to satisfy the type contract for r and the positivity contract for h, while precisely violating only the positivity contract for r.
-By pruning logically inconsistent combinations in advance, the solver ensures that only feasible and semantically valid test cases are generated.
-For our analysis of code generation, we then use these precisely targeted test cases to evaluate how different prompting strategies guide an LLM to produce implementations that properly enforce the intended contracts.
-In summary, current benchmarks overlook contract violations, thereby inflating the perception of correctness.
-Addressing this limitation, we propose the PACT framework.
-PACT extends HumanEval+ and MBPP+ to generate test cases focused on contract violations and introduces novel metrics to quantify contract adherence.
+Current code generation benchmarks (Chen et al., 2021; Austin et al., 2021; Liu et al., 2023) primarily measure functional correctness on well-formed inputs, typically through pass@k (Kulal et al., 2019; Chen et al., 2021).
+In this setting, test cases are curated to satisfy (often implicit) input conditions, and generated programs are judged by input-output correspondence on these pre-filtered inputs.
+This evaluation is appropriate for verifying functional behavior within the intended input domain, but it leaves a critical blind spot: a program can look correct under pass@k while failing to reject ill-formed inputs that violate validity constraints.
+Considering real-world software, failing to rule out such inputs and accepting invalid input can silently propagate erroneous states and undermine reliability assumptions.
+A natural way to formalize this boundary is through contracts (Meyer, 1992).
+In this work, we focus on assertion-level contracts—preconditions and input validity constraints that are realized as explicit assertion checks in a code (e.g., type, shape, range constraints).
+Importantly, many widely used benchmarks implicitly rely on such contracts during dataset constructions: ill-formed inputs are filtered out, leaving only well-formed tests for evaluation (Liu et al., 2023).
+As a result, it is possible for a generated solution to pass functional tests while still accepting inputs that a developer expects the implementation to reject.
+Figure 1 illustrates this mismatch.
+We introduce ContractEval, a benchmark for evaluating whether generated programs implement contract checks and reject ill-formed inputs.
+ContractEval is built on HumanEval+ and MBPP+ (Liu et al., 2023) by augmenting each task with contract-violation (ill-formed) tests derived from reference assertions.
+We define rejection as raising AssertionError.
+Because available reference assertions are not guaranteed to be complete, ContractEval includes a contract-quality filtering step to remove misaligned contracts and evaluates contract satisfaction with respect to the aligned assertions,
+A key technical challenge is generating contract-violation tests that target specific assertions.
+When multiple assertions interact, violating a chosen subset while satisfying the rest requires a sophisticated procedure to reflect dependencies among contract assertions.
+ContractEval addresses this via a neuro-symbolic pipeline that combines an LLM and an SMT solver (Barrett and Tinelli, 2018).
+An LLM converts assertions into SMT-compatible constraints, and an SMT solver enumerates satisfiable violation combinations to generate concrete inputs that violate selected assertions while satisfying others.
+This yields contract-targeted counterexamples that support fine-grained contract evaluation.
+Using ContractEval, we find that standard prompting frequently produces programs that pass well-formed tests but fail to reject contract-violating inputs.
+We further show that injecting a small number of contract-violation examples into the prompt substantially improves contract satisfaction while largely preserving functional correctness measured by pass@1.
+Figure 1: ContractEval provides test cases and metrics to evaluate whether the generated code satisfies contracts.
 2 Related work
-2.1 Automated Test and Code Generation
-Automated test and code generation have deep roots in traditional software engineering.
-Early test case generation was typically categorized into black-box methods like random and mutational fuzzing (Godefroid et al., 2005; Claessen & Hughes, 2000), white-box techniques such as symbolic execution (He et al., 2021; Cha et al., 2022; Godefroid et al., 2008), and grey-box approaches like coverage-guided fuzzing (Choi et al., 2019; Stephens et al., 2016; She et al., 2024; Qian et al., 2022).
-Similarly, early code generation relied on methods like probabilistic grammar-based frameworks (Bielik et al., 2016) and specialized language models (Feng et al., 2020).
-The advent of the Transformer architecture marked a paradigm shift, establishing LLMs as the dominant approach in both fields (Chen et al., 2021b).
-The development of modern Code LLMs now typically involves pre-training on code corpora, followed by instruction tuning (Ouyang et al., 2022) and refinement through more advanced techniques like reinforcement learning with execution feedback (Gehring et al., 2025).
-However, despite these advancements, the primary goal has remained the enhancement of functional correctness, typically measured by pass@k on benchmarks with well-formed inputs.
-2.2 Contracts
-The Design by Contract (DbC) paradigm (Meyer, 1992) argues that software reliability depends on explicitly stated preconditions, postconditions, and invariants, with run-time assertions as the enforcement mechanism.
-However, mainstream automated testing and existing test-generation benchmarks (Wang et al., 2025; Jain et al., 2025) often overlook latent defects from missing contract checks.
-While related work on failure-inducing test cases (Zhang et al., 2024; Zhong et al., 2025) is effective at causing general exceptions, this is not the same as verifying adherence to the precise semantics of a given contract.
-For instance, for a function that requires a list of positive numbers, a generic failure-inducing test might use None to cause a TypeError, but this does not verify the specific rule that all numbers must be positive.
-In contrast, a contract-aware test like [10, 20, -5] is designed to be caught precisely by an assertion checking for positive values.
+2.1 Code Generation Evaluation
+Most code generation benchmarks evaluate models by executing generated programs on unit tests and reporting pass@k (Wang et al., 2025; Jain et al., 2025), which captures functional input-output correctness on well-formed inputs (Chen et al., 2021; Hendrycks et al., 2021; Li et al., 2022).
+This evaluation standard is effective for verifying behavior within the intended input domain, but it typically does not measure whether a generated program reject ill-formed inputs that violate implicit validity constraints.
+Recent benchmark extensions (Chen et al., 2023; Liu et al., 2023) improve the coverage of functional behaviors by adding stronger or more diverse valid tests, yet their evaluation suites still primarily target well-formed inputs, leaving contract satisfaction largely unmeasured.
+2.2 Contract Validity with Assertions
+Design by Contract (DbC) treats preconditions, postconditions, and invariants as part of the specification, and emphasizes checking them in executable conditions such as assertion (Meyer, 1992).
+Contracts most commonly appear as assertion-level input validity checks that represent the intended input space.
+Prior work on robustness-oriented testing studies failure-inducing inputs that trigger generic exceptions (Zhang et al., 2024; Zhong et al., 2025), but a crashing input does not necessarily correspond to any particular contract assertion.
+For instance, assume a function that requires a list of positive numbers.
+An input None may trigger a TypeError regardless of the intended rule, “all numbers must be positive.” In contrast, contract evaluation requires tests that intentionally violate a chosen contract while remaining feasible with respect to others.
 This distinction highlights the need for systematic methods that can precisely target formal contract specifications rather than just triggering arbitrary errors.
-2.3 SMT Solver
-SMT (Satisfiability Modulo Theories) solvers are powerful engines for determining the satisfiability of complex logical formulas across various theories, such as arithmetic and strings (Barrett & Tinelli, 2018).
-They typically interface using SMT-LIB, a standardized formal language, with Z3 being a widely adopted implementation in automated testing (de Moura & Bjørner, 2008).
-Traditionally, techniques like symbolic execution have employed SMT solvers to find inputs that satisfy path constraints, focusing on functional coverage over well-formed inputs.
-In this paradigm, constraints are used as admission checks to filter out invalid data rather than as explicit targets for evaluation.
-In contrast, our framework leverages SMT solvers to systematically generate ill-formed inputs that precisely violate formalized contracts, allowing us to rigorously test for robustness.
-A detailed example of how we formulate contracts is provided in Appendix D.
-3 The Need for Contract-Aware Evaluation
-3.1 Why Contracts Matter: Blind Spots in Current Benchmarks
-Recent studies on code and test case generation (Korraprolu et al., 2025; Sung et al., 2025; Li et al., 2022) have dominantly relied on pass@k to evaluate functional correctness.
-While effective, these metrics, along with approaches that target failure-inducing inputs (Peng et al., 2018), share a common limitation: they primarily operate within the bounds of a program’s legal input space Dyck et al.
-(2023).
-This practice has led to a critical blind spot in popular benchmarks such as HumanEval+ and MBPP+.
-By design, these benchmarks explicitly filter out and discard any test case that violates a program’s pre-conditions (Austin et al., 2021).
-Consequently, the evaluation process certifies that a program behaves correctly on well-formed inputs, but it fails entirely to assess the program’s robustness against ill-formed ones.
-This results in an incomplete and often inflated assessment of code quality, praising solutions that may be superficially correct but are fundamentally fragile.
-This is where the software contracts become essential.
-As explained in Section 1, contracts are the rules that define the boundary of valid behavior.
-They are not optional safeguards but a core component of trustworthy software that specifies how a program must identify and reject invalid data.
-By ignoring contract adherence, existing benchmarks overlook a crucial dimension of software reliability.
-3.2 Contract-Based Test Paradigm
-Merely enlarging a single pool of valid inputs cannot reveal whether a model understands the boundary of the specification it needs to satisfy.
-The main focus, therefore, lies in constructing contract-violating test cases, which systematically explore the extent to which models enforce contract rules.
-A contract-violating test case is an input that violates one or more predicates from a set of contracts while satisfying the remaining specification.
-The reference implementation is augmented with run-time assertions for every predicate in the contract set, and a candidate program passes such a test only when it raises an error consistent with this augmented reference.
-Introducing contract-violating inputs uncovers false negatives that purely functional tests overlook and provides a rigorous measure of whether a program properly enforces contractual rules.
-3.3 Task Setup
-Each benchmark task consists of a natural language description, a set of contracts, and a functional implementation.
-We use HumanEval+ and MBPP+, where contract predicates are provided as assertion literals outside the prompt and reference code.
-Our task concerns contract-violating test generation and contract-aware code generation.
-For test generation, we automatically construct a compact set of contract-violating test cases that target specific contract rules and remain feasible with respect to the remaining rules for each benchmark task.
-These tests are used for evaluating whether LLM-generated code appropriately follows contracts.
-For contract-aware code generation, we generate code under two prompt conditions, where the first condition is a contract specification (CS) prompt, which includes the functional description and a natural language paraphrase of the contracts.
-The second condition is an example-augmented specification (EAS) prompt, which is augmented with the contract-violating test cases.
-Figure 2: Running example of PACT with an Example-Augmented Specification (EAS) prompt, which integrates Contract Specification (CS) and contract-violating test cases (CVTs) to enforce contract-aware code generation.
-4 Methodology
-We propose PACT, a program assessment and contract-adherence evaluation framework designed to systematically evaluate and enhance the ability of LLMs to generate contract-compliant code.
-This framework consists of two main stages.
-First, we generate contract-violating test cases to rigorously assess whether LLM-generated code snippets enforce both functional specifications and explicit contracts.
-Second is the systematic analysis of code generation, where we use these test cases under different prompting conditions to evaluate a model’s contract awareness in detail.
-Unlike prior approaches that rely solely on functionality-based evaluation, PACT extends the evaluation paradigm with contract-violation tests, enabling a more precise and reliable analysis of contract adherence.
-4.1 Contract-Violating Test-Case Generation
-Direct LLM-based generation is inadequate for contract-aware test case construction.
-It must consider all subsets of the contract set when producing violating inputs.
-If a task has $n$ contracts, the number of non-empty violation combinations is $2^{n}-1$.
-Direct prompting often misses required combinations and yields contradictory inputs that violate unintended contracts and fails to ensure feasibility under the specification.
-LLMs also lack a built-in mechanism to verify that exactly the targeted contracts are violated while all others hold.
-We design an SMT-based approach to make these checks accurate and efficient.
-Our generation procedure is a two-step pipeline for contract-aware test case generation.
-First, an inference model translates natural language contracts into an Algebraic Data Type (ADT) program.
-This ADT format is the critical component: it provides a rigid, formal schema for the complex and often nested constraints found in contracts.
-This ensures that the subsequently generated rules for an SMT solver are syntactically valid and semantically precise.
-Second, given the ADT, the SMT solver constructs contract violation test cases (CVTs) by breaking one or more specified contracts while ensuring that all remaining contracts are satisfied.
-The solver validates satisfiability for each candidate combination and extracts concrete models—specific assignments of values that satisfy the constraints— to instantiate inputs for every valid case.
-We provide a running example of this procedure in Appendix D.
-4.2 Contract Metrics for Test Case
-Unlike standard test cases for functional correctness, test cases for contracts should be the negative samples, violating the contracts and triggering the corresponding assertions.
-As the concept of contract-violating test cases is different from the standard, we design metrics to analyze whether the generated test cases appropriately correspond to input contracts.
-Let $A=\{a_{1},\ldots,a_{n}\}$ be the set of all contract assertions and $T=\{t_{1},\ldots,t_{m}\}$ be the generated test cases.
-$F_{t}\subseteq A$ is the set of violated assertions when executing a test case $t\in T$.
-Finally, $T_{neg}\subseteq T$ represents the set of test cases that successfully violated at least one contract.
-Assert Violation Coverage (AVC)
-AVC quantifies the coverage of assert statements that are successfully violated by all the test cases.
-$AVC=\frac{|\{a_{i}\mid\exists t\in T_{neg}:a_{i}\in F_{t}\}|}{n}.$
-The value of 1.0 for AVC ensures that all contract assertions are captured by the test cases. Lower values expose the unexplored regions of the contracts from the test cases.
-Target Specificity (TS)
-TS evaluates the precision of each test case by measuring how accurately it violates its intended set of target contracts.
-To formalize this, we first define $V\subseteq A$ as the set of contracts that a given test case $t\in T_{neg}$ is intended to violate.
-We then measure the alignment between this intended set $V_{t}$ and the actually violated set $F_{t}$ using the Jaccard Index.
-The final TS score is the average of these individual precision scores across all negative test cases:
-$TS=\frac{1}{|T_{neg}|}\sum_{t\in T_{neg}}\frac{|F_{t}\cap V_{t}|}{|F_{t}\cup V_{t}|}.$
-A score of 1.0 indicates perfect precision, where every test case violates exactly the set of contracts it was designed to.
-Lower scores reveal a discrepancy, indicating that test cases either failed to trigger their intended violations or caused unintended, collateral violations.
-4.3 Contract-Aware Code Generation
-This section details our methodology for systematically evaluating and enhancing the ability of LLMs to generate code snippets that robustly enforce contracts.
-To achieve a fine-grained understanding of a model’s contract awareness, we investigate its performance under two distinct prompting strategies.
-First, we establish a baseline condition, Contract Specification (CS).
-For this strategy, we use a powerful Commercial LLM to naturally integrate the contract rules from the HumanEval+ and MBPP+ datasets into the main text of the prompt.
-This creates a comprehensive prompt that describes both the functional goal and its contractual constraints in natural language.
-Second, we introduce an enhanced condition, which we term Example-Augmented Specification (EAS).
-This strategy builds upon the CS prompt by augmenting it with a single, precisely targeted contract- violating test cases for each described contract.
-This provides a concrete example of what constitutes a violation, intended to guide the model toward more robust code generation.
-A concrete example of this EAS prompt is shown in Figure 2.
-To measure the impact of these prompting strategies, we assess the generated code snippets using a comprehensive suite of metrics for both functional correctness and contract adherence.
-Functional correctness is measured using the standard pass@k metric over a set of valid test cases.
-Contract adherence is evaluated from two perspectives: runtime enforcement using test cases and static alignment of the generated code snippets with the ground-truth contracts, as detailed in Section 4.4.
-4.4 Contract Metric for Code Generation
-For the evaluation of the generated code snippets, we employ pass@k to measure functional correctness on valid inputs and use AVC to measure the correctness of the generated contract assertions.
-We also design two additional metrics to more precisely evaluate contract-aware code generation.
-Let $A=\{a_{1},\ldots,a_{n}\}$ be the set of ground-truth contract assertions, and let $\hat{A}=\{\hat{a}_{1},\ldots,\hat{a}_{m}\}$ be the set of assertions extracted from the LLM-generated code.
-Let $M\subseteq A\times\hat{A}$ be the set of pairs $(a_{i},\hat{a}_{j})$ where the ground-truth contract $a_{i}$ and the generated assertion $\hat{a}_{j}$ are determined to be semantically equivalent.
-Assertion Alignment Recall (AAR)
-AAR measures the model’s ability to implement all required contracts without omission.
-It is the proportion of ground-truth contracts that are successfully covered by at least one assertion in the generated code, functioning as a recall metric.
-$AAR=\frac{|\{a_{i}\in A\mid\exists\hat{a}_{j}\in\hat{A}:(a_{i},\hat{a}_{j})\in M\}|}{n}.$
-A high AAR score indicates that an output code ensures that all required contract specifications are generated.
-Assertion Alignment Precision (AAP)
-AAP measures the accuracy of the generated assertions, penalizing irrelevant or hallucinated ones.
-It is the proportion of generated assertions that correspond to a valid ground-truth contract, functioning as a precision metric.
-$AAP=\frac{|\{\hat{a}_{j}\in\hat{A}\mid\exists a_{i}\in A:(a_{i},\hat{a}_{j})\in M\}|}{m}.$
-A high AAP score indicates that the code does not contain assertions of unnecessary or incorrect checks.
+ContractEval targets this assertion-level notion of contracts and evaluates whether the generated programs reject ill-formed inputs in the intended way, using reference assertions available in HumanEval+ and MBPP+ as a starting point.
+2.3 SMT Solver for Test Generation
+SMT (Satisfiability Modulo Theories) solvers decide the satisfiability of logical formulas under background theories (e.g., arithmetic, strings), and are widely used in program analysis and constraint-based test generation (Barrett and Tinelli, 2018; de Moura and Bjørner, 2008).
+Classic uses include symbolic execution and path-constraint solving to synthesize inputs that satisfy feasibility constraints (Cadar et al., 2008).
+ContractEval uses SMT solving in a complementary way: we convert contract assertions into solver constraints and query for test cases that violate a selected subset of the assertions while satisfying the rest.
+This enables contract-targeted ill-formed test generation, which is difficult to achieve reliably with LLM-only generation.
+3 Observation on Contract Satisfaction
+HumanEval+ and MBPP+ provide assertion-level contracts as a dedicated dataset field, but their default test suites exclude contract-violating inputs by construction.
+As a result, standard evaluation certifies functional behavior only on well-formed inputs and leaves contract satisfaction untested.
+We expose this blind spot by executing generated implementations on contract-violating tests (CVTs) and measuring whether the implementation rejects them.
+We compare a standard benchmark prompt and a prompt augmented with negative examples.
+Figure 2 shows that the generated code rejects none of the CVTs under both prompts, indicating that negative examples do not elicit reliable contract checks.
+We then evaluate a contract-aware prompt that states the intended input constraints, and the same models reject substantially more CVTs under this prompt.
+This suggests that commonly used benchmark prompts often fail to generate programs that follow the provided input constraints in practice, unless they are surfaced explicitly.
+Figure 2: Evaluates contract satisfaction (contract sat.) of generated code. Standard and negative-example prompts yields 0% while a contract prompt achieves 50% performance on contract satisfaction.
+These findings motivate contract-aware evaluation on benchmarks with explicit contracts, where CVTs are executed rather than filtered out.
+Because this evaluation assumes that the provided contracts accurately reflect the task specification, we assess contract–specification alignment and filter out misaligned cases.
+3.1 HumanEval+ and MBPP+ for Contract-Aware Analysis
+HumanEval+ and MBPP+ annotate each task with an explicit contract field alongside the natural language description and reference solution.
+This design enables analyses that are difficult to perform in standard code generation benchmarks.
+First, we can directly test contract satisfaction by executing implementations on ill-formed inputs that violate the contract and checking whether they are rejected.
+Second, we can validate the contracts themselves by comparing their semantics with the task specification and determining whether they are specification-aligned.
+However, standard evaluation pipelines are not equipped to leverage these opportunities, as they are designed to use only well-formed inputs.
+By addressing this gap, we utilize HumanEval+ and MBPP+ as a natural foundation.
+3.2 Contract–Specification Alignment Check
+Contract satisfaction evaluation assumes that each contract captures mandatory input constraints stated or implied by the task description.
+If a contract is irrelevant, overly restrictive, or internally contradictory, contract-based evaluation becomes noisy.
+We therefore assess contract–specification alignment using an LLM judge and exclude tasks with misaligned contracts before downstream analyses.
+Figure 3: Overview of ContractEval construction.
+4 ContractEval Benchmark
+We introduce ContractEval, a code-generation benchmark with contract-violating tests (CVTs) to evaluate whether generated program reject invalid inputs specified by assertion-level contracts.
+Figure 3 illustrates the detail of our neuro-symbolic construction pipeline using both an LLM and SMT solver.
+Each ContractEval task consists of: (i) the original benchmark prompt and well-formed functional tests for pass@k, (ii) contract-aware query reconstructions by injecting intended contract information to original queries.
+and (iii) verified CVTs derived from reference assertions for contract satisfaction.
+We further introduce new evaluation metrics to quantify contract satisfaction under contract-violating tests.
+4.1 Contract-Violating Test Construction
+CVT Synthesis Pipeline.
+We propose a neuro-symbolic pipeline that combines the semantic translation capability of LLMs with the satisfiability guarantees of SMT solvers.
+The pipeline produces CVTs that violate a chosen subset of contracts while remaining feasible with respect to the remaining contracts within two stages.
+The first stage converts natural language level contract assertions into SMT-LIB constraints defined over an Algebraic Data Type (ADT).
+The second stage enumerates combinations of contracts, checks satisfiability for each combination, and synthesizes concrete inputs from satisfying combinations.
+Formalization with a Canonical ADT Template.
+Programs may accept inputs that combine primitive and structured values, including nested containers with mixed element types.
+Such inputs are difficult to encode directly in SMT because most solvers assume statically typed signatures and require explicit constructors for structured data.
+We address this challenge with a canonical value ADT embedded in an SMT-LIB template, as detailed in Appendix F.
+The ADT provides a unified representation for integers, floats, strings, booleans, and nested containers, allowing contracts to be encoded as SMT constraints over a single representation using shared constructors and auxiliary predicates.
+This uniform formalization supports systematic construction of constraint sets for different combinations of contract satisfaction and violation, which enables query generation for all non-empty violation combinations.
+An LLM then populates the template by converting the reference contracts into ADT-level constraints and instantiating the corresponding template slots.
+The instantiated template specifies the contract constraints for the solver on top of the fixed ADT representation.
+This template-driven translation yields a syntactically valid SMT-LIB instance and preserves the semantics of contracts required for the solver to recognize.
+Combinatorial Synthesis via SMT Solving.
+The pipeline performs combinatorial synthesis to generate test cases that cover distinct and logically feasible violation scenarios implied by the contract set.
+For a task with $n$ contracts, we consider all $2^{n}-1$ non-empty subsets as violation targets.
+For each target subset, we construct an SMT query that negates the targeted contract predicates and enforces satisfaction of the remaining predicates.
+The solver checks satisfiability and discards UNSAT targets that conflict with the remaining constraints.
+For each SAT query, the solver returns a ADT-level model assignment, which constitutes a CVT in the encoded space.
+We decode this assignment into an executable test input by converting ADT constructors into their corresponding runtime values and recursively rebuilding nested containers.
+The resulting CVTs are precisely targeted and logically consistent.
+Feasibility Filtering for CVTs.
+Test cases synthesized by the SMT solver are logically consistent with respect to the contract predicates.
+Logical consistency alone does not ensure that an input is suitable for evaluating contract satisfaction.
+A valid CVT must be functionally feasible while still violating the intended contract.
+The same input should execute successfully when contract checks are removed, and it should fail with the intended assertion when those checks are enabled.
+Inputs that crash with generic runtime exceptions, such as TypeError or IndexError, even without contracts, do not isolate contract enforcement and instead reflect functional invalidity.
+We filter synthesized CVTs by executing each input on the reference implementation under two configurations: (i) with assertions removed to verify functional feasibility, and (ii) with assertions intact to confirm the targeted contract violation.
+This filtering reduces the task set from 419 to 364 because some tasks admit no input that is both executable without contracts and contract violating with assertions enabled.
+The resulting tasks ensure that ContractEval measures contract enforcement rather than robustness to generic errors.
+4.2 Contract-Aware Query Reconstruction
+Each task in ContractEval includes an original natural-language query and a set of reference contract assertions.
+We use a commercial LLM to reconstruct a contract-aware query by interpreting the intent of each assertion and rewriting the original query so that the corresponding input constraints are stated explicitly in natural language.
+This step standardizes how contract information is presented to code generation models and supports controlled comparisons by ensuring that evaluations focus on the model’s ability to satisfy explicitly stated constraints.
+4.3 ContractEval Metrics
+ContractEval reports CVT quality metrics, assertion violation capture (AVC) and target specificity (TS) to validate that synthesized tests are clause-targeted rather than generic crashing inputs.
+ContractEval also reports functional correctness on well-formed inputs using pass@k, and contract satisfaction of generated programs using contract satisfaction rate (CSR) below.
+CVT Quality Metrics (AVC, TS).
+Let $A=\{a_{1},\ldots,a_{n}\}$ be the set of contract assertions and $T=\{t_{1},\ldots,t_{m}\}$ be the generated test cases.
+For a test case $t\in T$, let $F_{t}\subseteq A$ denote the set of assertions violated when executing $t$.
+AVC measures the fraction of generated test cases that successfully trigger at least one assertion violation:
+$\mathrm{AVC}=\frac{|\{t\mid t\in T\wedge|F_{t}|>0\}|}{|T|}.$
+A higher AVC indicates that the generator more consistently produces CVTs that elicit at least one assertion failure.
+A lower AVC indicates that a larger portion of generated tests do not trigger any assertion failure.
+TS evaluates how each CVT matches its intended contract targets.
+Let $T_{\mathrm{neg}}=\{t\in T\mid|F_{t}|>0\}$ be the set of tests that trigger at least one assertion violation.
+For each $t\in T_{\mathrm{neg}}$, let $V_{t}\subseteq A$ denote the set of assertions that $t$ is intended to violate.
+TS measures the Jaccard similarity between the intended target set $V_{t}$ and the actually violated set $F_{t}$ and averages it over $T_{\mathrm{neg}}$:
+$\mathrm{TS}=\frac{1}{|T_{\mathrm{neg}}|}\sum_{t\in T_{\mathrm{neg}}}\frac{|F_{t}\cap V_{t}|}{|F_{t}\cup V_{t}|}.$
+A TS of $1$ indicates that each test violates exactly its intended target set. A smaller TS indicates that a test misses some intended violations, triggers unintended violations, or both.
+Contract Satisfaction Rate (CSR)
+To evaluate contract enforcement against contract-violating inputs, we use verified CVTs obtained from our filtering process.
+Let $T_{\mathrm{ver}}=\{t_{1},\ldots,t_{k}\}$ denote the verified CVT set.
+For a generated code snippet $C$, let $T_{\mathrm{rej}}\subseteq T_{\mathrm{ver}}$ be the subset of inputs that $C$ rejects by raising an assertion failure or an explicit rejection.
+CSR quantifies the fraction of verified CVTs that are rejected by $C$:
+$\mathrm{CSR}=\frac{|\{t\mid t\in T_{\mathrm{ver}}\wedge t\in T_{\mathrm{rej}}\}|}{|T_{\mathrm{ver}}|}.$
 5 Experimental Settings
 5.1 Datasets
-Our study utilizes HumanEval+ and MBPP+ benchmarks (Liu et al., 2023), which are enriched with assertion-level contracts.
-However, these benchmarks have a critical limitation for evaluating code robustness: the contract specifications are neither included in the model prompts, nor are contract-violating inputs included in the official test suites.
-This evaluation setup exclusively tests for functional correctness on well-formed inputs, leading to an inflated perception of code quality.
-We construct a supplementary dataset containing contract-violating test cases, as described in Section 4.1.
-5.2 Candidate Models
-We utilize o4-mini as our primary test case generator.
-For our code generation experiments, we evaluate a set of open-source models including gemma-3-12B-it (gemma-3), Deepseek-R1-Distill-Qwen-14B (DeepSeek-R1), Qwen3-14B (Qwen-3), and Phi-4-reasoning-plus (Phi-4-plus).
+We evaluate on ContractEval built on HumanEval+ and MBPP+ (Liu et al., 2023).
+For each task, we use the original well-formed tests to measure functional correctness and a set of verified CVTs, which we constructed, to measure contract satisfaction.
+We run all experiments on the filtered task set with specification-aligned contracts and feasible CVTs.
+5.2 Prompting Strategies
+We evaluate three prompt formats: standard, contract specification (CS), and example-augmented specification (EAS).
+The standard prompt uses the original natural-language query.
+CS uses the contract-aware rewrite from Section 4.2.
+EAS appends a small set of verified contract-violating tests (CVTs; Section 4.1) to CS as negative examples.
+We also evaluate Chain-of-Thought (CoT) (Wei et al., 2022) and a multi-turn prompting variant.
+5.3 Evaluation Metrics
+As described in Section 4.3 We evaluate contract-violating test generation using AVC and TS.
+For code generation, we report pass@k on well-formed tests and CSR on verified CVTs, and additionally use CodeBLEU Ren et al.
+(2020) and an LLM-as-judge score to assess the syntactic and semantic quality of the generated assertion checks.
+5.4 Evaluated Models
+We employ o4-mini as the primary engine for contract-violating test case generation, contract-aware query reconstruction, and as an LLM-as-judge.
+For our code generation experiments, we evaluate a set of open-source models including gemma-3-12B-it (gemma-3), Deepseek-R1-Distill-Qwen-14B (DeepSeek-R1), Qwen3-14B (Qwen-3), Phi-4-reasoning (Phi-4), and Phi-4-reasoning-plus (Phi-4-plus).
+We provide detailed implementation settings in Appendix A for completeness and reproducibility.
 6 Empirical Studies
-Our empirical study is structured over three research questions (RQs) designed to evaluate the PACT framework from multiple perspectives.
-The evaluation is conducted on HumanEval+ and MBPP+.
-We begin by assessing the quality and precision of the CVTs generated by PACT (RQ1).
-Next, we investigate whether providing these concrete test cases is more effective for eliciting contract-awareness than using abstract natural language descriptions alone (RQ2).
-Finally, we analyze the resulting trade-off between the enforced contract adherence and the functional correctness of the generated code snippets (RQ3).
-Table 1: Evaluation results of the test case generation on HumanEval+ and MBPP+.
-Benchmark Method AVC ($\uparrow$}) TS ($\uparrow$}) AVG ($\uparrow$})
-HumanEval+ o4-mini 97.14% 75.60% 86.37%
-o4-mini + SMT Solver 95.53% 85.81% 90.67%
-MBPP+ o4-mini 94.67% 69.11% 81.89%
-o4-mini + SMT Solver 93.50% 84.54% 89.02%
-6.1 RQ1: How Effective is PACT in Generating High-Quality Contract- Violating Test Cases?
-Our first research question investigates the effectiveness of our proposed framework, PACT, in generating high-quality CVTs.
-A naive baseline approach, which we refer to as o4-mini in Table 1, uses an LLM o4-mini for direct test case generation from given contracts.
-This method is often inadequate, as it disregards dependencies among contracts and produces logically inconsistent violations, making it difficult to precisely verify individual contract predicates.
-In contrast, PACT (o4-mini+ SMT Solver) employs a more robust two-stage approach.
-First, an LLM converts contracts into rules for an SMT solver.
-Subsequently, an SMT solver uses these rules to generate test cases.
-This process guarantees the generation of logically sound test cases that are precisely targeted to violate a specific subset of contracts while adhering to the rest.
-The empirical results in Table 1 validate the superiority of PACT.
-While both methods achieve high AVC scores, PACT significantly outperforms the direct generation in TS, achieving over 10%p increases on both HumanEval+ and MBPP+.
-This demonstrates that our two-stage approach successfully gives test cases precisely for intended contracts.
-The baseline’s slightly higher AVC is an expected outcome, as direct generation often triggers a wide range of assertion errors indiscriminately, artificially inflating coverage.
-PACT, on the other hand, focuses on valid, targeted violations.
-While PACT generates valid rules, providing a precise test for the input contracts, there are still errors.
-A case-specific analysis of minor errors that occur even with well-formed rules is detailed in Appendix A.
-Table 2: Evaluation results of contract adherence in code generation on HumanEval+.
-Model Mode pass@1 ($\uparrow$}) AVC ($\uparrow$}) AAR ($\uparrow$}) AAP ($\uparrow$}) AVG ($\uparrow$})
-gemma-3 CS 84.41% 24.85% 11.41% 14.04% 32.79%
-EAS 82.94% 91.02% 28.07% 27.77% 57.45%
-DeepSeek CS 73.78% 44.12% 15.65% 16.97% 37.63%
-EAS 71.77% 79.29% 27.62% 28.01% 51.67%
-Qwen3 CS 78.92% 28.04% 13.17% 22.55% 35.67%
-EAS 77.83% 87.81% 31.53% 36.09% 58.31%
-Phi-4-plus CS 72.23% 52.91% 18.78% 21.09% 41.25%
-EAS 67.08% 69.50% 21.33% 20.06% 44.49%
-Table 3: Evaluation results of contract adherence in code generation on MBPP+.
-Model Mode pass@1 ($\uparrow$}) AVC ($\uparrow$}) AAR ($\uparrow$}) AAP ($\uparrow$}) AVG ($\uparrow$})
-gemma-3 CS 78.56% 57.99% 17.50% 17.93% 41.49%
-EAS 78.60% 95.57% 32.29% 31.82% 59.57%
-DeepSeek CS 62.53% 64.20% 17.59% 17.57% 40.47%
-EAS 60.15% 86.70% 28.23% 27.94% 45.47%
-Qwen3 CS 72.41% 70.86% 21.09% 22.99% 46.84%
-EAS 72.63% 94.85% 31.54% 32.30% 57.83%
-Phi-4-plus CS 64.89% 67.33% 24.26% 24.65% 45.28%
-EAS 63.76% 74.88% 29.20% 28.95% 49.20%
-6.2 RQ2: Are Concrete Test Cases More Effective than Abstract Descriptions for Eliciting Contract-Awareness?
-We compared two prompting methods across various models in Tables 2 and 3.
-CS, our baseline, provides only an abstract natural language description of the contracts.
-In contrast, EAS augments the CVTs, generated from Section 4.1, to the CS prompt.
-The results demonstrate that for all of our base models, EAS is a more powerful method for eliciting contract-aware code generation.
-EAS achieves 33.7%p, 11.3%p, and 9.4%p improvements for AVC, AAR, and AAP, respectively, over CS.
-For models such as Qwen3, DeepSeek-R1, and Phi-4-plus, switching from CS to EAS leads to a dramatic improvement across all contract adherence metrics.
-This indicates that concrete examples of CVTs provide a clear and unambiguous signal that forces the model to move beyond purely functional logic.
-The increase in AAR and AAP indicates that the models are correctly generating test cases required for assertion checks, while the relatively huge improvements in AVC show that the resulting codes are more robust against a wider range of invalid inputs.
-Concrete examples of contract violations provide a clear, unambiguous signal that forces the model to move beyond purely functional logic and implement the explicit enforcement of contracts—the predefined agreements on how ill-formed inputs must be rejected.
-For concrete examples of the code snippets and their corresponding LLM-generated test cases, please refer to Appendix B.
-However, this heightened focus on contract enforcement introduces a notable trade-off, as the model may generate more complex code that impacts its performance on purely functional correctness metrics.
-This trade-off is analyzed in further detail in the following section RQ3.
-6.3 RQ3: How Does Enforcing Contract Adherence Impact the Functional Correctness of Generated Code?
-While the RQ2 established that augmenting prompts with CVTs significantly enhances a model’s contract adherence, this section investigates the resulting trade-off with respect to its functional correctness.
-Our analysis reveals that a heightened focus on contract enforcement often comes at the cost of a measurable decline in performance on standard functional correctness benchmarks.
-This trade-off can be attributed to a shift in the model’s optimization focus.
-When provided with only a functional description, the LLM’s sole objective is to produce code snippets that pass functionality-focused tests.
-However, when the prompt is augmented with CVTs, the model must simultaneously satisfy two competing objectives: generating functionally correct logic and enforcing the specified contractual preconditions.
-This dual objective increases the complexity of the code generation task.
-The model is compelled to allocate part of its reasoning capacity to interpreting and implementing the contract rules, which can lead to subtle errors or oversights in the core functional logic.
-A concrete example of this is presented in Appendix C.
-Consequently, while the resulting code snippets are more robust and secure against invalid inputs, they may exhibit a lower pass@1 rate on test suites composed entirely of valid, well-formed inputs.
-In conclusion, our findings reveal an inherent tension between contract adherence and functional correctness in LLM-based code generation.
-The pursuit of robustness through contract enforcement comes at a tangible cost to functionality.
-This trade-off highlights the critical need for advanced training paradigms—such as the reinforcement learning with multi-objective rewards proposed by PACT—capable of optimizing for both objectives simultaneously.
+Our empirical study answers three research questions about ContractEval.
+We evaluate the validity of ContractEval by checking that reference assertions are aligned with the task specification and that CVTs precisely target selected contract assertions while remaining feasible under the others.
+Using ContractEval, we quantify a blind spot of standard code generation evaluation: models fail to reject CVTs.
+Finally, we study whether prompting strategies beyond our default setting can improve the trade-off between functional correctness and contract satisfaction.
+6.1 Validating ContractEval: Contract Alignments and CVTs
+ContractEval evaluates whether generated programs check assertion-level contracts—preconditions and validity constraints encoded as reference assertions—by testing them on CVTs (ill-formed inputs).
+This evaluation is valuable when (i) the reference assertions align with the task specification and (ii) the ill-formed tests precisely target intended contract clauses, rather than triggering arbitrary failures.
+Table 1: Comparison of contract-violating test cases generated by o4-mini and by ContractEval (ours).
+Method AVC ($\uparrow$}) TS ($\uparrow$}) Average ($\uparrow$})
+o4-mini 97.58% 72.69% 85.13%
+ContractEval (ours) 94.11% 84.92% 89.52%
+While ContractEval brings assertions from HumanEval+ and MBPP+, they are not designed as a contract-evaluation benchmark and do not provide a guarantee that the assertions are semantically aligned with the task description.
+We perform a contract-quality check using o4-mini as an LLM judge that assesses whether the assertions reflect input constraints stated or implied by the task specification.
+Across 542 tasks that include reference assertions, 88.04% are judged specification-aligned.
+We filtered out programs with misaligned assertions for accurate downstream contract evaluation, as explained in Section 4.1.
+This filtering removes noise so that subsequent analyses reflect the model capability more precisely.
+Table 2: Functional correctness and contract satisfaction in code generation.
+Adding contract descriptions (CS) helps, and adding a few ill-formed examples (EAS) yields large gains in contract satisfaction.
+Model Instruction Functional Contract Satisfaction
+pass@1 ($\uparrow$) CSR ($\uparrow$) CodeBLEU ($\uparrow$) LLM-as-judge ($\uparrow$)
+DeepSeek-R1 Standard 80.81% 0.00% 0.00% 0.00%
+CS 73.29% 27.24% 34.29% 53.05%
+EAS 70.87% 52.66% 54.51% 81.49%
+gemma-3 Standard 81.76% 0.00% 0.00% 0.00%
+CS 74.75% 22.69% 29.97% 41.74%
+EAS 69.56% 48.54% 48.88% 72.04%
+Phi-4 Standard 75.39% 0.00% 0.00% 0.00%
+CS 71.29% 40.71% 46.66% 69.33%
+EAS 72.99% 50.73% 50.86% 75.63%
+Phi-4-plus Standard 75.97% 0.00% 0.00% 0.00%
+CS 70.83% 38.24% 44.38% 65.31%
+EAS 73.21% 50.94% 51.04% 75.98%
+Qwen-3 Standard 74.79% 0.00% 0.00% 0.00%
+CS 70.62% 24.29% 35.05% 49.71%
+EAS 69.65% 51.83% 53.30% 77.27%
+We validate that ContractEval’s ill-formed tests accurately probe intended contracts.
+We compare o4-mini as a direct LLM baseline that generates ill-formed tests to our neuro-symbolic synthesis pipeline, ContractEval.
+Table 1 shows that both approaches achieve high AVC, but our approach yields substantially higher TS, indicating that ContractEval produces more precise and targeted CVTs for contract combinations.
+While ContractEval shows high AVC performance, its score is not 100% accurate, indicating that additional filtering is mandatory for precise evaluation.
+A CVT should be ill-formed with respect to the contract while still being feasible as an execution input.
+Otherwise, a test may trigger unrelated errors that do not reflect contract satisfaction.
+Thus, we apply a feasibility filtering detailed in Section 4.1.
+We take intersecting tasks that pass both filtering, which results in 364 tasks remaining.
+These filtering processes remove tests that are ill-typed or semantically irrelevant and ensures that our evaluation reflects intended contract checking behavior.
+We provide a case analysis of detailed errors in Appendix C.
+6.2 Contract-Aware Code Generation
+Recall that we use ContractEval to evaluate how well LLMs generate contract-checking assertions that reject CVTs in addition to functional logic.
+Specifically, we measure functional correctness with pass@1 on well-formed tests and contract satisfaction on CVTs.
+We compare 1) a standard prompt with task description, 2) a CS prompt with both task and contract description, and 3) a EAS prompt of CS with a small set of CVT examples.
+From Table 2, it is straightforward that a standard prompt for all base LLMs yields zero performance on contract evaluation.
+While the standard prompt achieves strong functional accuracy with 75–82% pass@1 scores, at the same time, it yields 0.00 scores on all contract satisfaction metrics.
+This shows that passing conventional well-formed tests does not imply that the generated program implements the intended input-validity checks.
+ContractEval makes this blind spot measurable: without explicit steering with contract information, models tend to implement functional logic while silently accepting invalid inputs, leading to missing or incomplete assertion logic.
+We first examine the CSR performance of CS and EAS as CSR accurately evaluates whether the generated contracts successfully raises AssertionErrors for CVTs.
+Adding contract constraints in natural language descriptions, we achieve 22–41%p increases in CSR, indicating that models translate abstract constraints into assertion checks.
+However, EAS yields the largest gains in CSR by accomplishing 10–30%p over CS.
+On average, EAS achieves 50.94%p CSR improvement and maintains 92% of a pass@1 score, compared with the standard prompt.
+While CSR is a precise and verifiable metric to check contract satisfaction of LLM-generated codes, we also provide CodeBLEU and LLM-as-judge scores for syntactic and semantic evaluation.
+Note that the standard prompt still yields 0.00 scores for these two metrics.
+EAS performs the best for these two metrics, gaining 13.65%p and 20.65%p performance improvements over CS.
+Even among functionally correct generations, contract satisfaction remains far from perfect.
+For EAS, when we restrict evaluation to tasks where the generated solution achieves pass@1=1=1, the mean CSR stays in the 0.51–0.59 range across models (Appendix H).
+Thus, functional correctness and contract satisfaction are related but not equivalent objectives, reinforcing the need for a benchmark that evaluates both.
+6.3 How do Prompt Strategies Affect Contract-Aware Code Generation?
+After establishing that EAS is effective for eliciting contract-aware behavior, we analyze whether additional prompting techniques improve the trade-off between functional correctness and contract satisfaction.
+We consider (i) Chain-of-Thought (CoT) prompting and (ii) a refinement-style multi-turn strategy that first generates a solution with functional logic and then refines the code with contract information.
+Table 3 summarizes the impact of these techniques across our base LLMs.
+While CoT slightly improves pass@1 by 1.2%p for gemma-3 and Qwen-3, it frequently degrades both pass@1 and CSR among other LLMs.
+More critically, CoT catastrophically fails on Phi-4 and Phi-4-plus.
+These models collapse to 0–4% on both metrics.
+This suggests that CoT is not appropriate in our setting, where the output must remain executable and simultaneously demonstrate the intended rejection behavior on CVTs.
+In contrast, multi-turn is stable across models and improves functional correctness over EAS in most cases.
+In most cases, EAS achieves superior CSR performance to that of multi-turn, however, multi-turn achieves better pass@1 performance, improving 1.8%p over EAS on average.
+For instance, Phi-4-plus increases pass@1 from 73.21% to 77.42%, and DeepSeek-R1 increases from 70.87% to 72.96%.
+However, this gain is often accompanied by a modest drop in CSR by 2–7%p, indicating that refinement tends to prioritize preserving functional behavior over implementing every contract check implied by CVTs.
+Table 3: Comparison of additional prompting strategies. Multi-turn recovers pass@1 while maintaining competitive contract satisfaction.
+Model Prompting pass@1 ($\uparrow$}) CSR ($\uparrow$})
+DeepSeek-R1 EAS 70.87% 52.66%
++ CoT 61.11% 35.00%
++ Multi-Turn 72.96% 45.44%
+gemma-3 EAS 69.56% 48.54%
++ CoT 70.46% 50.37%
++ Multi-Turn 71.34% 49.26%
+Phi-4 EAS 72.99% 50.73%
++ CoT 3.66% 1.91%
++ Multi-Turn 71.88% 46.58%
+Phi-4-plus EAS 73.21% 50.94%
++ CoT 0.00% 0.00%
++ Multi-Turn 77.42% 48.52%
+Qwen-3 EAS 69.65% 51.83%
++ CoT 71.13% 45.24%
++ Multi-Turn 71.70% 49.23%
+Overall, EAS remains the strongest single-turn strategy for contract satisfaction, while a multi-turn strategy is the most reliable way to regain pass@1.
+Yet, none of the tested prompting variants achieves both high pass@1 and near-perfect CSR, highlighting and motivating future methods evaluated under ContractEval.
 7 Conclusion
-We introduce PACT, the first framework to redefine code and test case correctness by evaluating adherence to task specifications through both functionality and contract-based behavior.
-While prior benchmarks assess only pass@k on well-formed inputs, PACT introduces a comprehensive paradigm with dual test suites—one for functionality and one for contract violations—along with specific metrics to analyze contract awareness and uncover latent defects, enabling a more principled evaluation of code robustness.
-Our empirical evaluation demonstrates the effectiveness of PACT across multiple dimensions.
-We first show that PACT’s SMT-solver-based test case generation method ensures more accurate CVTs than direct generation with over 10%p better performance.
-Furthermore, our results reveal that augmenting prompts with these CVTs is a highly effective strategy for generating robust and contract-aware code, achieving 18.13%p increase in contract-specific metrics and 12.8%p in total average.
-Our empirical studies on various LLMs demonstrate the effectiveness of PACT, achieving approximately, but also uncovers a critical trade-off between this enhanced contract adherence and a model’s performance on functional correctness.
-These findings confirm that PACT provides a more complete and realistic assessment of an LLM’s contract-aware code generation capabilities, moving beyond the limitations of existing benchmarks.
-8 Future work directions
-While this work focuses on evaluating contract adherence, a natural next step is to actively improve it through advanced training methodologies.
-A promising direction is to leverage the novel metrics introduced in our PACT framework as direct training signals.
-The novel metrics introduced in our PACT framework, such as the runtime metric Assert Violation Coverage (AVC) and the static metrics Assertion Alignment Recall (AAR) and Precision (AAP), are particularly well-suited for this purpose and could be integrated into a multi-objective reward function for RL.
-This approach could enable models to learn to navigate the trade-off between functional correctness and contract adherence more effectively, optimizing for both objectives simultaneously.
-Appendix A Case Study: Logical Contradictions in Direct LLM Test Case Generation
+ContractEval reframes code generation evaluation by testing not only functional correctness on well-formed inputs, but also whether generated program reject CVTs via intended assertions.
+Built on HumanEval+ and MBPP+, ContractEval augments each task with contract-targeted CVTs.
+We propose a neuro-symbolic pipeline to construct precise CVTs where an LLM converts contract assertions into SMT-recognizable constraints and an SMT solver enumerates satisfiable violation combinations.
+ContractEval further introduces metrics to quantify (i) the precision of generated CVTs and (ii) contract satisfaction of generated programs.
+Empirically, we find that standard prompts for code generation yields 0% for contract satisfaction, failing to reject CVTs.
+This reveals a blind spot in conventional benchmarks.
+By contrast, EAS substantially improves contract satisfaction by 50.94% while preserving 92% of pass@1, compared to the standard prompt.
+Overall, ContractEval provides a benchmark and methodology for evaluating assertion-level contract checks as an essential axis of robustness in LLM-based code generation.
+8 Limitations
+ContractEval is designed to make assertion-level contract satisfaction measurable and comparable, but several practical constraints define its current scope and point to natural extensions.
+Scope of contracts.
+ContractEval derives contracts from reference assertions, which may be incomplete with respect to all validity constraints that developers intend.
+We mitigate misalignment by filtering tasks whose assertions are judged inconsistent with the natural-language specification, but ContractEval does not claim to exhaustively capture every admissible/invalid input boundary for each task.
+Coverage of invalid inputs.
+CVTs are clause-targeted and feasibility-checked, but they still represent a finite set of contract-violation patterns.
+They may not cover all semantically invalid inputs (e.g., distributional shifts, adversarial edge cases, or domain-specific constraints not expressed as assertions).
+Future versions could combine ContractEval’s targeted CVTs with complementary robustness tests (e.g., fuzzing-style perturbations) while preserving interpretability.
+Cost and evaluation dependencies.
+SMT enumeration and feasibility filtering introduce computational overhead compared to purely LLM-based test generation, and any auxiliary model-based judgments (e.g., contract–spec alignment checks or semantic scoring) may inherit bias from the judge model.
+Reducing reliance on model-based judgments and improving efficiency are practical directions for scaling ContractEval.
+Appendix A Implementation Detail
+[…]
+Appendix B Hidden Contracts and Contract-Violating Tests
+Natural-language programming tasks rarely specify only functional behavior.
+They also imply a boundary of admissible inputs—type restrictions, format requirements, range constraints, and other validity rules that developers take for granted.
+In this paper, we call these validity rules contracts and refer to contract-violating tests (CVTs) as inputs that violate one or more contract clauses (i.e., ill-formed inputs with respect to the intended input space).
+Figure 4: Hidden contracts and contract-violating tests (CVTs). A typical LLM solution can pass prior benchmarks that only test well-formed inputs while silently accepting contract-violating inputs.
+Figure 4 illustrates why CVTs matter for code generation evaluation.
+The task asks for has_dup(nums) that returns whether an input sequence contains duplicates.
+A typical LLM-generated solution correctly implements the core logic using a set, and it will pass standard unit tests that only cover well-formed inputs.
+However, the task description carries a hidden contract that nums is a list of numbers.
+Without explicit input checks, the same program may silently accept contract-violating inputs (e.g., a string or a mixed-type list) and still return a plausible output.
+This creates a blind spot: pass@k can certify a solution as “correct” without ever testing whether it rejects contract-violating inputs.
+ContractEval makes this missing axis measurable by augmenting each task with CVTs derived from assertion-level contracts.
+Concretely, CVTs are designed so that a contract-checking implementation should raise an error (e.g., by triggering an assertion) instead of producing a normal output.
+Beyond evaluation, the same CVT examples can be reused as a lightweight steering signal: including a small number of CVTs in the prompt encourages models to generate explicit contract checks (assertions) alongside the functional logic.
+Appendix C Case Study: Logical Contradictions in Direct LLM Test Case Generation
 HumanEval
@@
 This highlights a fundamental weakness of the approach, as the LLM tends to generate simplistic contract-violation test cases that fail to respect the logical relationships among contracts.
+HumanEval/113 def odd_count(lst): assert type(lst) == list, "invalid inputs" # \$_CONTRACT_\$ assert all(isinstance(s, str) for s in lst), "invalid inputs" # \$_CONTRACT_\$ assert all(s.isdigit() for s in lst), "invalid inputs" # \$_CONTRACT_\$ ans, template = [], "the number of odd elements in the string i of the input." for s in lst: odd_cnt = len(list(filter(lambda ch: int(ch) % 2 == 1, s))) ans.append(template.replace("i", str(odd_cnt))) return ans """ Contract List: assert_0: assert type(lst) == list, "invalid inputs assert_1: assert all(isinstance(s, str) for s in lst), "invalid inputs assert_2: assert all(s.isdigit() for s in lst), "invalid inputs """"
+Figure 5: Code and contracts for HumanEval.
 Mbpp
@@
 This highlights a fundamental weakness of the approach, as the LLM tends to generate simplistic contract-violation test cases that fail to respect the logical relationships among contracts.
-HumanEval/113 def odd_count(lst): assert type(lst) == list, "invalid inputs" # \$_CONTRACT_\$ assert all(isinstance(s, str) for s in lst), "invalid inputs" # \$_CONTRACT_\$ assert all(s.isdigit() for s in lst), "invalid inputs" # \$_CONTRACT_\$ ans, template = [], "the number of odd elements in the string i of the input." for s in lst: odd_cnt = len(list(filter(lambda ch: int(ch) % 2 == 1, s))) ans.append(template.replace("i", str(odd_cnt))) return ans """ Contract List: assert_0: assert type(lst) == list, "invalid inputs assert_1: assert all(isinstance(s, str) for s in lst), "invalid inputs assert_2: assert all(s.isdigit() for s in lst), "invalid inputs """"
-Figure 3: Code and contracts for HumanEval.
 MBPP/731 import math def lateralsurface_cone(r,h): assert isinstance(r, (int, float)), "invalid inputs" # \$_CONTRACT_\$ assert isinstance(h, (int, float)), "invalid inputs" # \$_CONTRACT_\$ assert r > 0, "invalid inputs" # \$_CONTRACT_\$ assert h > 0, "invalid inputs" # \$_CONTRACT_\$ l = math.sqrt(r * r + h * h) return math.pi * r * l """ Contract List: assert_0: assert isinstance(r, (int, float)), "invalid inputs assert_1: assert isinstance(h, (int, float)), "invalid inputs assert_2: assert r > 0, "invalid inputs assert_3: assert h > 0, "invalid inputs """
 Figure 6: Code and contracts for MBPP.
-Appendix B Case Study: Contract Enforcement under Different Prompting Conditions
+Appendix D Case Study: Contract Enforcement under Different Prompting Conditions
 We present a direct comparison of code snippets that the DeepSeek model generated for the MBPP/11 task under two prompting conditions.
 This comparison illustrates why providing CVTCs is more effective than relying on natural language descriptions alone.
+MBPP/11 """ Write a python function that removes the first and last occurrence of a given character from the string; the first input must be a non-empty string and the second input must be a string of length one.
+""" def remove_Occ(s, char): assert len(s) > 0, "First input must be a non-empty string." assert len(char) == 1, "Second input must be a string of length one." first = s.find(char) last = s.rfind(char) if first == -1: return s return s[:first] + s[first+1:last] + s[last+1:]
+Figure 7: Code generated by DeepSeek with the contract specification (CS) prompt.
 Figure 7 shows the code snippets produced when the model received only the natural language prompt.
@@
 MBPP/11 """ Write a python function that removes the first and last occurrence of a given character from the string; the first input must be a non-empty string and the second input must be a string of length one.
-""" def remove_Occ(s, char): assert len(s) > 0, "First input must be a non-empty string." assert len(char) == 1, "Second input must be a string of length one." first = s.find(char) last = s.rfind(char) if first == -1: return s return s[:first] + s[first+1:last] + s[last+1:]
-Figure 5: Code generated by DeepSeek with the contract specification (CS) prompt.
-MBPP/11 """ Write a python function that removes the first and last occurrence of a given character from the string; the first input must be a non-empty string and the second input must be a string of length one.
 # Contract Test Cases: >>> remove_Occ(2.0, 3.0) "AssertionError: invalid input" >>> remove_Occ(5.0, 4.0) "AssertionError: invalid input" >>> remove_Occ(’’, 2.0) "AssertionError: invalid input" >>> remove_Occ(2.0, ’’) "AssertionError: invalid input" """ def remove_Occ(s, char): assert isinstance(s, str) and len(s) > 0, "invalid input" assert isinstance(char, str) and len(char) == 1, "invalid input" first_occ = s.find(char) last_occ = s.rfind(char) if first_occ == -1: return s return s[:first_occ] + s[first_occ+1:last_occ] + s[last_occ+1:]
 Figure 8: Code generated by DeepSeek with the example-augmented specification (EAS) prompt.
-Appendix C Case Study: Contract Adherence vs. Functional Correctness
+Appendix E Case Study: Contract Adherence vs. Functional Correctness
 A case study from the HumanEval/142 task provides a concrete example of the trade-off between contract adherence and functional correctness.
 The figures 911 compare the ground-truth implementation (golden label), the code generated from a contract natural language prompt (CS), and the code generated from a prompt augmented with contract-violating test cases (CVTs).
+HumanEval/142 def sum_squares(lst): assert type(lst) == list, "invalid inputs" # \$_CONTRACT_\$ assert all(type(x) == int for x in lst), "invalid inputs" # \$_CONTRACT_\$ ans = 0 for i, num in enumerate(lst): if i % 3 == 0: ans += num ** 2 elif i % 4 == 0: ans += num ** 3 else: ans += num return ans
+Figure 9: The ground-truth implementation, correct in both functionality and contracts.
 Figure 10 displays the code generated from the NL-only prompt, which describes both functional requirements and contractual constraints.
@@
 This outcome suggests the model prioritized achieving functional correctness—even developing a logic for handling multiples of 12—at the complete expense of implementing the required robustness checks, leaving the code snippets vulnerable to invalid inputs.
+HumanEval/142 def sum_squares(lst): total = 0 for i in range(len(lst)): num = lst[i] if i % 3 == 0 and i % 4 == 0: total += num ** 2 elif i % 3 == 0: total += num ** 2 elif i % 4 == 0: total += num ** 3 else: total += num return total
+Figure 10: A functionally correct implementation that fails to enforce contracts.
 In contrast, Figure 11 shows the code generated when the prompt was augmented with CVTCs.
@@
 This underscores the inherent tension between these two competing objectives in LLM-based code generation.
-HumanEval/142 def sum_squares(lst): assert type(lst) == list, "invalid inputs" # \$_CONTRACT_\$ assert all(type(x) == int for x in lst), "invalid inputs" # \$_CONTRACT_\$ ans = 0 for i, num in enumerate(lst): if i % 3 == 0: ans += num ** 2 elif i % 4 == 0: ans += num ** 3 else: ans += num return ans
-Figure 7: The ground-truth implementation, correct in both functionality and contracts.
-HumanEval/142 def sum_squares(lst): total = 0 for i in range(len(lst)): num = lst[i] if i % 3 == 0 and i % 4 == 0: total += num ** 2 elif i % 3 == 0: total += num ** 2 elif i % 4 == 0: total += num ** 3 else: total += num return total
-Figure 8: A functionally correct implementation that fails to enforce contracts.
 HumanEval/142 def sum_squares(lst): assert isinstance(lst, list), "AssertionError: invalid input" for elem in lst: assert isinstance(elem, int), "AssertionError: invalid input" total = 0 for index, num in enumerate(lst): if index % 3 == 0 and index % 4 != 0: total += num ** 2 elif index % 4 == 0 and index % 3 != 0: total += num ** 3 else: total += num return total
 Figure 11: A robust implementation that enforces contracts but fails on functionality.
-Appendix D Formalizing Contracts into SMT-LIB: Template and a Case Study
-This section details the structure of the SMT-LIB template used by our PACT framework.
+Appendix F Formalizing Contracts into SMT-LIB: Template and a Case Study
+This section details the structure of the SMT-LIB template used by the construction of ContractEval.
 SMT-LIB is a standardized, text-based language used to interface with SMT solvers.
@@
 Our framework leverages this language to translate nuanced, natural language contracts into a formal representation that can be reasoned about with logical precision.
+The SMT-LIB template ADT_BASE_TEMPLATE = """ (set-logic ALL) ; ==== CANONICAL PYTHON-LIKE ADT (DO NOT MODIFY) ==== (declare-datatypes ((Value 0)) ( ((IntVal (ival Int)) (FloatVal (fval Real)) (StrVal (sval String)) (BoolVal (bval Bool)) (Nil) (Cons (head Value) (tail Value))) )) ; === ADD HELPER FUNCTIONS HERE === <<HELPER_FUNCTIONS>> ; === Inputs === <<INPUT>> ; === BASIC STRUCTURE === <<BASIC_STRUCTURE>> ; === Contract predicates === <<CONTRACT_DEFS>> ; === COMBINATION === <<COMBINATION>> (check-sat) (get-model) """
+Figure 12: The SMT-LIB template used for formalizing contracts.
 Figure 12 shows the base template we designed for this purpose.
@@
 The SMT solver then attempts to find a concrete model that satisfies this exact combination of constraints.
+HumanEval/11 from typing import List def string_xor(a: str, b: str) -> str: assert isinstance(a, str) and isinstance(b, str), "invalid inputs" # \$_CONTRACT_\$ assert len(a) == len(b), "invalid inputs" # \$_CONTRACT_\$ assert set(a).issubset({"0", "1"}) and set(b).issubset({"0", "1"}), "invalid inputs" # \$_CONTRACT_\$ return "".join(str(int(a[i]) ^ int(b[i])) for i in range(len(a)))
+Figure 13: The ground-truth implementation in HumanEval/11
 Figure 13 shows the ground-truth Python implementation for the HumanEval/11 task, which requires a function that takes two binary strings of equal length.
 The SMT-LIB formalization of these requirements is shown in Figure 14.
-The three assert statements in the Python code directly correspond to the three formal contracts defined in SMT-LIB:
-  • C0 verifies that both inputs are strings, corresponding to the assertion assert isinstance(a, str) and isinstance(b, str).
+The three assert statements in the Python code directly correspond to the three formal contracts defined in SMT-LIB: C0 verifies that both inputs are strings, corresponding to the assertion assert isinstance(a, str) and isinstance(b, str).
 C1 ensures their lengths are equal, corresponding to assert len(a) == len(b).
 C2 checks that they are valid binary strings using a custom isBinaryString helper function, corresponding to assert set(a).issubset("0", "1") and set(b).issubset("0", "1").
+The SMT-LIB template ADT_BASE_TEMPLATE = """ (set-logic ALL) ; ==== CANONICAL PYTHON-LIKE ADT (DO NOT MODIFY) ==== (declare-datatypes ((Value 0)) ( ((IntVal (ival Int)) (FloatVal (fval Real)) (StrVal (sval String)) (BoolVal (bval Bool)) (Nil) (Cons (head Value) (tail Value))) )) ; === ADD HELPER FUNCTIONS HERE === (define-fun Safe_Sval ((x Value)) String (ite (is-StrVal x) (sval x) "")) (define-fun isBinaryString $($s Value)) Bool (and (is-StrVal s) (str.in.re (Safe_Sval s) (re.* (re.union (str.to.re "0") (str.to.re "1")))))) ; === Inputs === (declare-const a Value) (declare-const b Value) ; === BASIC STRUCTURE === ; === Contract predicates === (define-fun C0 () Bool (and (is-StrVal a) (is-StrVal b))) (define-fun C1 () Bool (= (str.len (Safe_Sval a)) (str.len (Safe_Sval b)))) (define-fun C2 () Bool (and (isBinaryString a) (isBinaryString b))) ; === COMBINATION === (assert (not C0) (assert (not C1) (assert (not C2) (check-sat) (get-model) """
+Figure 14: An example of the SMT-LIB template populated for HumanEval/11
 The COMBINATION block determines the goal of the test case generation.
@@
 The specific instance in the figure 14, for example, asserts the negation of all three contracts to generate a test case that violates every precondition simultaneously.
-The SMT-LIB template ADT_BASE_TEMPLATE = """ (set-logic ALL) ; ==== CANONICAL PYTHON-LIKE ADT (DO NOT MODIFY) ==== (declare-datatypes ((Value 0)) ( ((IntVal (ival Int)) (FloatVal (fval Real)) (StrVal (sval String)) (BoolVal (bval Bool)) (Nil) (Cons (head Value) (tail Value))) )) ; === ADD HELPER FUNCTIONS HERE === <<HELPER_FUNCTIONS>> ; === Inputs === <<INPUT>> ; === BASIC STRUCTURE === <<BASIC_STRUCTURE>> ; === Contract predicates === <<CONTRACT_DEFS>> ; === COMBINATION === <<COMBINATION>> (check-sat) (get-model) """
-Figure 10: The SMT-LIB template used for formalizing contracts.
-HumanEval/11 from typing import List def string_xor(a: str, b: str) -> str: assert isinstance(a, str) and isinstance(b, str), "invalid inputs" # \$_CONTRACT_\$ assert len(a) == len(b), "invalid inputs" # \$_CONTRACT_\$ assert set(a).issubset({"0", "1"}) and set(b).issubset({"0", "1"}), "invalid inputs" # \$_CONTRACT_\$ return "".join(str(int(a[i]) ^ int(b[i])) for i in range(len(a)))
-Figure 11: The ground-truth implementation in HumanEval/11
-The SMT-LIB template ADT_BASE_TEMPLATE = """ (set-logic ALL) ; ==== CANONICAL PYTHON-LIKE ADT (DO NOT MODIFY) ==== (declare-datatypes ((Value 0)) ( ((IntVal (ival Int)) (FloatVal (fval Real)) (StrVal (sval String)) (BoolVal (bval Bool)) (Nil) (Cons (head Value) (tail Value))) )) ; === ADD HELPER FUNCTIONS HERE === (define-fun Safe_Sval ((x Value)) String (ite (is-StrVal x) (sval x) "")) (define-fun isBinaryString $($s Value)) Bool (and (is-StrVal s) (str.in.re (Safe_Sval s) (re.* (re.union (str.to.re "0") (str.to.re "1")))))) ; === Inputs === (declare-const a Value) (declare-const b Value) ; === BASIC STRUCTURE === ; === Contract predicates === (define-fun C0 () Bool (and (is-StrVal a) (is-StrVal b))) (define-fun C1 () Bool (= (str.len (Safe_Sval a)) (str.len (Safe_Sval b)))) (define-fun C2 () Bool (and (isBinaryString a) (isBinaryString b))) ; === COMBINATION === (assert (not C0) (assert (not C1) (assert (not C2) (check-sat) (get-model) """
-Figure 12: An example of the SMT-LIB template populated for HumanEval/11
+Appendix G Full Result of Prompting Strategies
+Table 4 reports the full metrics for the prompting variants discussed in Section 6.3.
+Beyond pass@1 and CSR, we include CodeBLEU and LLM-as-judge to capture syntactic and semantic alignment of the generated contract checks with the reference.
+Overall, CoT prompting is high-variance and often harmful: it can substantially degrade both functional correctness and contract satisfaction, and it catastrophically collapses for Phi-4 and Phi-4-plus (near-zero across all metrics).
+In contrast, the multi-turn strategy is markedly more stable.
+It typically recovers pass@1 relative to the single-turn EAS baseline (e.g., up to +4.2%p for Phi-4-plus), while keeping CSR competitive (within roughly 0–7%p of EAS across models).
+The auxiliary metrics largely track CSR: when CSR drops under a prompting variant, CodeBLEU and LLM-as-judge tend to drop as well, suggesting that performance changes are driven by missing or altered contract checks rather than superficial formatting differences.
+Table 4: Comparison of additional prompting strategies. Multi-turn recovers pass@1 while maintaining competitive contract satisfaction to the standard EAS prompt.
+Model Instruction Functional Contract Satisfaction
+pass@1 ($\uparrow$) CSR ($\uparrow$) CodeBLEU ($\uparrow$) LLM-as-judge ($\uparrow$)
+DeepSeek-R1 EAS 70.87% 52.66% 54.51% 81.49%
++ CoT 61.11% 35.00% 40.35% 62.46%
++Multi-Turn 72.96% 45.44% 45.75% 65.34%
+gemma-3 EAS 69.56% 48.54% 48.88% 72.04%
++ CoT 70.46% 50.37% 50.68% 71.23%
++Multi-Turn 71.34% 49.26% 49.62% 72.45%
+Phi-4 EAS 72.99% 50.73% 50.86% 75.63%
++ CoT 3.66% 1.91% 8.14% 13.28%
++Multi-Turn 71.88% 46.58% 48.44% 73.42%
+Phi-4-plus EAS 73.21% 50.94% 51.04% 75.98%
++ CoT 0.00% 0.00% 2.92% 4.96%
++Multi-Turn 77.42% 48.52% 49.87% 75.01%
+Qwen-3 EAS 69.65% 51.83% 53.30% 77.27%
++ CoT 71.13% 45.24% 50.01% 75.62%
++Multi-Turn 71.70% 49.23% 50.97% 76.98%
+Appendix H Contract Evaluation on Functionally Correct Programs
+Through ContractEval, our experiments show that prompting LLMs with contract information (e.g., CS and EAS) substantially improves contract satisfaction on ill-formed tests.
+However, contract satisfaction is only meaningful when the generated program is also functionally correct on well-formed inputs.
+If a program already fails on valid inputs, then its behavior on ill-formed inputs is ambiguous: it may “reject” an input simply because it crashes (e.g., due to unrelated bugs), rather than because it implements the intended contract checks.
+To disentangle true contract behavior from incidental failures, we analyze contract satisfaction conditioned on functional correctness in both Figure 15 and Table 5.
+Figure 15: Contract satisfaction depending on the performance of functaional correctness.
+Let $f_{i}\in[0,1]$ denote the functional test pass rate of a generated program for task $i$ (measured on well-formed tests), and let $c_{i}\in[0,1]$ denote its contract satisfaction (measured by CSR on ContractEval’s ill-formed tests).
+For a functionality threshold $t$, we consider the subset $\mathcal{S}_{t}=\{i\mid f_{i}\geq t\}$ and report: (i) the conditional mean contract satisfaction $\mathbb{E}[c\mid f\geq t]$, and (ii) the probability of perfect contract satisfaction $\Pr(c=1\mid f\geq t)$.
+Intuitively, increasing $t$ filters to progressively more functionally reliable programs, and reveals whether contract satisfaction improves as functional correctness improves.
+Across models, we observe that contract satisfaction does not reliably track functional correctness.
+After removing the most functionally incorrect outputs (e.g., $t\geq 0.25$), $\mathbb{E}[c\mid f\geq t]$ becomes largely flat: making a program more functionally correct does not automatically make it more contract-satisfying.
+Even among fully functionally correct programs ($f=1$), a substantial fraction still fails to reject contract-violating inputs.
+For example, under our contract-aware prompting setting, only about half of functionally correct programs achieve perfect contract satisfaction ($\Pr(c=1\mid f=1)$ is roughly in the 0.51–0.59 range across models), indicating that many solutions that “solve the task” still omit essential validity checks.
+This conditional analysis supports the central premise of ContractEval: functional correctness on well-formed inputs is insufficient to characterize robustness.
+Contract satisfaction remains a distinct and non-trivial dimension of code quality, and ContractEval’s ill-formed tests expose failures that are invisible under pass@k-style evaluation.
+Table 5: Contract satisfaction evaluation on functionally correct programs.
+Model Instruction Correct Functions Contract Satisfaction
+Ratio ($\uparrow$) CSR ($\uparrow$) CodeBLEU ($\uparrow$) LLM-as-judge ($\uparrow$)
+DeepSeek-R1 CS 63.19% 32.89% 35.96% 55.38%
+EAS 59.34% 56.75% 55.01% 83.01%
+gemma-3 CS 64.84% 26.53% 31.98% 44.61%
+EAS 57.14% 54.29% 54.20% 79.99%
+Phi-4 CS 61.26% 52.01% 52.25% 78.34%
+EAS 61.81% 61.78% 56.33% 84.21%
+Phi-4-plus CS 59.89% 49.58% 48.11% 71.89%
+EAS 62.91% 61.82% 55.31% 83.31%
+Qwen-3 CS 60.99% 30.05% 39.93% 55.83%
+EAS 59.34% 58.76% 57.79% 85.18%